### v2

# Proficient Graph Neural Network Design by Accumulating Knowledge on Large Language Models

Abstract
High-level automation is increasingly critical in AI, driven by rapid advances in large language models (LLMs) and AI agents. However, LLMs, despite their general reasoning power, struggle significantly in specialized, data-sensitive tasks such as designing Graph Neural Networks (GNNs). This difficulty arises from (1) the inherent knowledge gaps in modeling the intricate, varying relationships between graph properties and suitable architectures and (2) the external noise from misleading descriptive inputs, often resulting in generic or even misleading model suggestions. Achieving proficiency in designing data-aware models—defined as the meta-level capability to systematically accumulate, interpret, and apply data-specific design knowledge—remains challenging for existing automated approaches, due to their inefficient construction and application of meta-knowledge. To achieve the meta-level proficiency, we propose DesiGNN, a knowledge-centered framework that systematically converts past model design experiences into structured, fine-grained knowledge priors well fitted to meta-learning with LLMs. To account for the inherent variability and external noise, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs. By constructing a solid meta-knowledge between unseen graph understanding and known effective architecture patterns, DesiGNN can deliver top-5.77% initial model proposals for unseen datasets within seconds, and achieve consistently superior performance with minimal search costs against baselines.
1 Introduction
Automation has become a defining trend in artificial intelligence, reshaping how complex tasks are executed at scale without human intervention or the requirement of deep expertise. From code generation [7, 9, 60] and agentic workflow orchestration [57, 30, 64] to the multi-task generalization of foundation models [29], this paradigm shift is transforming how expertise is embedded and operationalized in AI systems. Among various frontiers, the automatic design of machine learning models stands as one of the most fundamental and impactful challenges in this automation landscape. In the past, automated machine learning (AutoML) [44, 45, 48, 62] has over-emphasized model search specific to independent data and tasks from scratch, which does not conform to the cross-domain and cross-task automation trends. The recent breakthroughs in large language models (LLMs) [11, 4, 41] and AI agents [25, 19, 63] with strong academic backgrounds and contextual reasoning have further fueled expectations—could they autonomously generate tailored models for unseen tasks, thereby democratizing machine learning?
Inspired by LLMs’ general reasoning and generation capability, recent studies [40] have leveraged LLMs to replace human involvement beyond traditional abilities of AutoML, such as task understanding [43, 12] and problem formulation [49]. Despite democratizing complex processes of designing models, realizing highly data-aware automated model design, especially for data-sensitive domains like graph learning, remains elusive for LLMs. Unlike tasks in vision or language domains where LLMs excel in proposing generic solutions that could dominate all tasks and data properties, designing Graph Neural Networks (GNNs) [24, 59, 52, 10, 18, 42] for specific graph properties and graph learning tasks requires a nuanced understanding of how data topology influences model architecture [54, 47]. For example, the homophily nature behind GCN’s aggregation has catastrophic effects when propagating information over heterophily graphs, where two nodes with different labels are more likely to be linked. Unfortunately, current LLM-based methods often fully rely on LLMs’ inherent knowledge and reasoning ability, lacking fine-grained design knowledge about the relationship between graph-specific properties and effective GNN configurations, while also being sensitive to external input noise. As observed in Figure 1(a), the LLM agents adopted in recent methods [43, 12, 49] tend to suggest generally well-performing yet overly generic architectures, inherently falling short in tailoring designs to unseen graph structures. Worse still, as shown in Figure 1(b), they are vulnerable to “artificial hallucinations”—being easily misled by external, irrelevant user descriptions (e.g., being a citation network) or blindly enumerated data statistics that do not accurately reflect the data-aware challenges in model design, resulting in misleading knowledge association and ineffective designs.
(a) Inherent Knowledge Gap: LLMs propose overly generic GNNs when provided with purely semantic and property descriptions, unable to induce data-specific configurations.
(b) External Noise: LLMs misinterpret task similarity by weighing more on the citation networks category, causing “artificial hallucinations” and ineffective knowledge transfers.
Figure 1: Case studies illustrating inherent and external limitations of existing LLM-based methods.
Achieving proficiency in designing data-aware models thus demands far more than general LLM-based model design methods, which calls for a meta-level capability to systematically accumulate, interpret, and apply fine-grained, data-specific design knowledge. Humans, akin to training a new Ph.D. student, face substantial hurdles in accumulating this complex knowledge, often undergoing a lengthy learning process to design tailored models proficiently. Meta-based AutoML methods represent initial efforts toward this goal computationally, seeking transferable insights across tasks via learned task embeddings [54, 1, 26] or surrogate predictors [50, 51]. These methods, however, often rely on simple, predefined meta-features that may fail to generalize robustly when facing complex data diversity, thus bending the current automation trend. Extending this concept to graphs, AutoTransfer [5] has made the first promising strides by creating task-model repositories and embedding-based knowledge transfers to improve search efficiency. Nonetheless, it remains limited by predefined task embeddings and the rigid aggregation of prior tasks into random search, insufficiently capturing the fine-grained, graph-aware meta-knowledge documented in the vast graph learning literature [54, 47].
To overcome these limitations, we propose DesiGNN, a knowledge-centered and proficiency-oriented framework for the automated design of data-aware GNNs for unseen tasks. DesiGNN systematically converts past model design experiences into structured, fine-grained knowledge priors that are well fitted to meta-learning with LLMs, novelly bridging the long-standing gap between unseen graph understanding and known effective architecture patterns in graph application research. To address the observed external noise from descriptive inputs and the inherent variability in graph properties, DesiGNN constructs an empirically validated meta-level understanding about graph-GNN-performance from benchmarks and adaptively aligns key properties with graph-specific insights elicited from the graph learning literature, using LLMs as knowledge aligners to enhance the constructed meta-knowledge. This hybridization of empirical knowledge and LLM-guided reasoning empowers DesiGNN to deliver immediate, data-aware model proposals without any training, efficiently refine these architectures based on accumulated insights, and ultimately deliver superior performance with minimal computational overhead. Our contributions are as follows:
  • We empirically identify two key limitations in current LLM-based GNN design approaches: the detrimental external noise from descriptive inputs and the inherent gap in fine-grained design knowledge. Motivated by these insights, we propose a knowledge-centered framework that enhances meta-level proficiency in designing tailored, data-aware GNNs for unseen tasks by systematically accumulating, interpreting, and applying fine-grained meta-knowledge.
  • To account for the inherent variability of graph properties to GNN performance, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs, constructing a solid meta-knowledge between graph topological properties and high-performing GNN architectures. This enables DesiGNN to proficiently deliver immediate, data-aware GNN recommendations without any direct training on the unseen datasets.
  • Extensive experiments across 8 benchmarks and 3 out-of-distribution datasets demonstrate that DesiGNN delivers initial GNN proposals ranking in the top 5.77% of all possible architectures within seconds. Moreover, its knowledge-driven refinement strategy rapidly optimizes these initial suggestions, consistently outperforming automated methods in both performance and efficiency.
2 Related Work
2.1 Automated GNNs
Automated GNNs (AutoGNNs) [44, 15, 62] aim to automatically find an optimal GNN architecture $\theta$ for an unseen graph $G^{u}$ using a controller $\pi$ parameterized by $\alpha$, where the architecture $\theta$ with network weight $\omega$ can achieve best performance $\mathcal{H}(\theta,\omega;G^{u})$ on $G^{u}$. In AutoGNNs, the search spaces are categorized into intra-layer [14] and inter-layer [21] designs, with strategies including reinforcement learning [14], evolutionary algorithms [39], and differentiable search [46]. However, as introduced in Section 1, these approaches lack proficiency and demand substantial computational resources [33], as they optimize configurations $\theta$ without leveraging prior knowledge and require extensive sampling from $\pi_{\alpha}(\theta)$. Recently, meta-learning [13, 6] approaches have sought to overcome these limitations by transferring insights from past tasks through learned task embeddings [54, 5, 1] or surrogate predictors [50, 51, 28]. While these methods mark an initial step toward knowledge-driven AutoGNNs, they rely on coarse, predefined meta-features, often incapable of capturing the generalizable relationships between diverse data properties and effective model configurations.
NAS Benchmark for Designing GNNs. NAS-Bench-Graph [35] offers a standardized benchmark for GNN architecture search, comprising nine datasets and 26,206 unique architectures with associated performance records. Similarly, there are many other benchmarks created by pioneers in different domains [8]. While these benchmarks could provide valuable insights into model-performance relationships, existing methods underutilize their potential for knowledge-driven design due to the absence of a data dimension, which is the foundation of knowledge transfer. Therefore, only a few methods [54, 5] attempted to design feasible models for unseen tasks using such model-performance knowledge. Notably, they rely predominantly on semi-supervised or supervised techniques to fit the missing associations with the data, i.e., require direct training on the unseen data.
2.2 LLMs for Designing GNNs and the Limitations
Inspired by LLMs’ ability to solve optimization tasks [53], recent studies explore LLMs for automating model design [40, 61, 22, 60, 55, 58, 17], focusing on managing task formulation, architecture generation, and optimization. Auto2Graph [49] democratizes the usage of AutoGNNs by translating descriptive user inputs (e.g., Cora [37] is a citation network) into configuration specifications, while GPT4GNAS [43] and GHGNAS [12] iteratively refine designs through crafted prompts, focusing on different graph geneity. Despite democratizing the usage, they suffer from: (1) Over-reliance on external descriptive user inputs that result in poor alignment between graph properties and architecture designs; (2) Raw LLMs inherently lack deep expertise in fine-grained graph-GNN-performance relationships, limiting their ability to suggest data-aware models; (3) Refining model designs solely based on feedback from testing models, lacking actionable, knowledge-driven search strategies.
3 DesiGNN: Proficient Graph Neural Network Design
Figure 2: The illustration of our knowledge-centered DesiGNN pipeline for designing GNNs.
As discussed in Section 1, current LLM-based methods suffer from the lack of proficiency in designing data-aware GNNs, where inherent knowledge gaps and external noise are two major barriers. To formalize these challenges, we adopt a Bayesian perspective to represent the automated GNNs design as solving the posterior belief over a candidate architecture $\theta_{i}$ when facing an unseen graph $G^{u}$:
$\mathbb{P}(\theta_{i}|G^{u})=\frac{\mathbb{P}(G^{u}|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(G^{u})},$ (1)
where $\mathbb{P}(\theta_{i})$ is derived from a knowledge pool $\mathcal{K}$ accumulated from prior tasks $\{G^{i}\}$. However, it remains a non-trivial problem to solve pseudo-likelihood $\mathbb{P}(G^{u}|\theta_{i})$ without formally testing the model on unseen data—a common proficiency bottleneck in existing AutoML that leads to exhaustive search. This proficiency denotes the capability to systematically accumulate, interpret, and apply fine-grained, data-aware meta-knowledge relating graph properties to effective GNN architectures. From a Bayesian perspective, such capability can be conceptually represented as solving the posterior belief $\mathbb{P}(\theta_{i}|G^{u})$ with minimal or even without testing candidate models $\{\theta_{i}\}$ on $G^{u}$.
To enhance such meta-level proficiency without directly testing the unseen data, DesiGNN explicitly integrates empirically validated graph-property knowledge with adaptive LLM alignment to construct a trustworthy meta-level understanding $\mathcal{M}$ between graphs $\mathbf{G}$, GNNs $\mathbf{\Theta}$, and performances $\mathbf{P}$: $\mathcal{M}:(\mathbf{G},\mathbf{\Theta})\rightsquigarrow\mathbf{P}$ based on crucial insights from graph literature [54, 47]. Figure 2 illustrates our Bayesian-inspired pipeline: Graph Understanding, Knowledge Retrieval, and Model Suggestion and Refinement, each component systematically addressing external noise of descriptive data inputs and inherent meta-knowledge gaps, followed by how to strategically apply the constructed meta-knowledge in designing data-aware models. We present our pseudo-code in Algorithm 1 in Appendix A.1.
3.1 Connect Graph Topology to GNN Performance
(a) Property confidence varies across graphs.
(b) Descriptive inputs induce invalid task associations that deviate from empirical truth. DesiGNN corr.=0.52.
Figure 3: Empirical studies of two barriers: inherent property variability and external noise.
Performance-oriented modeling of discrete graph data is essential to quickly capture data priors from massive benchmarks to construct trustworthy meta-knowledge, a key challenge for LLMs due to their inability to naturally interpret graph-structured data [16]. As discussed in Section 2.2, existing LLM-based methods rely heavily on user-provided task descriptions, which fail to capture the key graph topology (e.g., heterophily) that frequently inspires specific GNN designs in the vast graph learning literature [54, 47]. Our studies in Figure 1(b),  3(b) and Appendix D.2 demonstrate that semantic or purely descriptive inputs (e.g., is a citation graph) frequently mislead LLMs, causing “artificial hallucination” and ineffective solutions. Such descriptive inputs introduce significant external noise, undermining the reliability of generic LLM suggestions when connecting graphs to model performance. Thus, it is necessary to establish an empirical foundation that captures fine-grained graph-GNN-performance mappings, enabling the construction of trustworthy meta-knowledge for data-aware model design.
Table 1: 16 graph topological properties
Clustering Betweenness Density Degree centrality
Closeness Degree Edge count Graph diameter
Shortest path Assortativity Eigenvector Dimensionality
Node count Diversity Connected Label homophily
To empirically study the missing meta-level understanding about graph-GNN-performance, we start with 16 graph properties $\{g_{k}\}_{k=1}^{16}$ from the literature (see Table 1 and Appendix A.2) to infer graph similarity from two aspects: 1) the statistical distance ranking $\mathcal{SR}(i,k)$ based on graph properties $g_{k}$, and 2) the empirical performance ranking $\mathcal{ER}(i)$ of transferred top-performing GNNs from other graphs $\{G^{j}\mid j\neq i\}$ to anchor graph $G^{i}$. The concrete definition is presented in Appendix A.2.2. To assess how reliably specific graph properties predict successful GNN design knowledge transfers, we introduce the concept of graph property confidence for $g_{k}$ on $G^{i}$ through Kendall’s $\tau$ Coefficient [23]:
$I(G^{i},g_{k}):=\text{KendallCorr}\big{(}\mathcal{SR}(i,k),\mathcal{ER}(i)\big{)},$ (2)
where $I(G^{i},g_{k})$ quantifies the correlation between the topological distance measured by $g_{k}$ and the empirical performance differences in transferring different top-performing GNNs from $\{G^{j}\mid j\neq i\}$ to $G^{i}$. A higher $I(G^{i},g_{k})$ indicates that $g_{k}$ is a stronger indicator of effective model design transfers, thus playing a critical role in meta-learning with LLMs to design data-aware models for unseen tasks.
[…]
Empirical Filtering as Trustworthy Foundation. Recognizing those empirical insights on external noise and inherent property variability, we first utilize empirical filtering to preliminarily identify reliably influential graph properties by averaging their predictive confidences $I(G^{i},g_{k})$ across benchmark graphs as $\bar{I}(g_{k})=\frac{1}{n}\sum_{i=1}^{n}I(G^{i},g_{k}).$ This process ensures that our property set built from the literature is also empirically grounded, reducing susceptibility to external noise and inherent property variability. Then, based on the empirical observation in Figure 6(b), we develop a filter $\mathcal{F}(G^{i},\bar{I})$ to correlate the modeling of discrete graph data with transferred model performance, selecting the Top-$N_{f}$ properties with the highest $\bar{I}(g_{k})$. This leads to the alternative formulation of Equation 1:
$\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))=\frac{\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(\mathcal{F}(G^{u}))}.$ (3)
This mechanism mitigates the typical LLM failure of “artificial hallucinations” often triggered by unprofessional graph modeling with negligible computational cost offline. To accumulate new knowledge, we design a self-evaluation bank $BE$, which stores measured graph properties and similarity rankings for new graphs on the fly, allowing $\mathcal{F}$ to scalably evolve over time by integrating new empirical evidence. Grounded in empirical transferability, this mechanism serves as a trustworthy foundation for capturing graph-to-graph ($G^{i}$ to $G^{j}$) meta-priors that underpin the correlations between graphs and high-performing GNN models, but still calling for an effective property variability handler.
3.2 Knowledge Retrieval Module
Proficiency in GNN design largely depends on capturing the meta-knowledge in the form of the posterior belief for unseen tasks. However, even after empirically filtering and establishing a general correlation between graph properties and model performance, solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ remains a resource-intensive accumulation of testing various $\theta_{i}$ on $G^{u}$. To overcome it, we novelly propose to utilize the known likelihood $\mathbb{P}(\mathcal{F}(G^{i})|\theta_{i})$ on the benchmark datasets to form a practical proxy for $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ on the unseen. Eliciting such fine-grained model design knowledge paves the way for efficient approximation of the posterior $\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))$ with accumulated prior knowledge. Finding such a practical proxy for the unseen task among the massive benchmark knowledge requires an unsupervised task similarity formulation that properly assesses the likelihood of successful model design knowledge transfer from $G^{i}$ to $G^{u}$, which has to reflect the performance outcomes $\mathcal{H}(\cdot,\cdot;G^{u})$.
Definition 1 (Task Similarity for Model Design Knowledge Transfer)
Given an unseen graph $G^{u}$ and two benchmark graphs $G^{i}$ and $G^{j}$, each associated with $N_{m}$ top-performing models $\{\theta^{*}_{m}\}_{m=1}^{N_{m}}\in\Theta$, an effective task similarity $\mathcal{S}(\cdot,\cdot)$ for transfering model design knowledge satisfies:
$\mathcal{S}(G^{u},G^{i})\geq\mathcal{S}(G^{u},G^{j})$ (4)
$\Rightarrow$ [($\mathbb{P}\Bigl{[}\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})\geq\mathcal{H}(\{\theta^{*}_{jm}\},\cdot,G^{u})\Bigr{]}\geq\delta.$
where ($\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$ denotes the performance of the transferred top-performing pattern $\{\theta^{*}_{im}\}$ from benchmark $G^{i}$ on the unseen $G^{u}$, and $\delta$ is a confidence threshold. This ensures that higher task similarity between $G^{u}$ and $G^{i}$ implies a larger likelihood of high-performing transfer from $G^{i}$ to $G^{u}$.
To design such performance-oriented task similarity between $G^{u}$ and $G^{i}$ unsupervisedly, we first apply the filter $\mathcal{F}(\cdot)$ (initialized with accumulated benchmark knowledge in Section 3.1) on $G^{u}$ and $G^{i}$ to establish an empirical foundation based on performance-oriented confidence $\bar{I}(g_{k})$. However, as discussed in Section 3.1, graph properties inherently exhibit varying levels of confidences across graphs, making it challenging to transfer knowledge reliably with standalone statistical properties. To account for this inherent variability, we further introduce the adaptive importance $w^{u}_{k}$ for each property $g_{k}$ specifically tailored to the unseen graph $G^{u}$ and formulate our unsupervised task similarity as:
$\mathcal{S}(G^{u},G^{i})=\frac{1}{N_{f}}\sum_{k=1}^{N_{f}}w^{u}_{k}\cdot\frac{\bar{I}(g_{k})}{1+\hat{d_{k}^{ui}}},$ (5)
where $N_{f}$ is the number of influential graph properties selected via empirical filtering and $\bar{I}(g_{k})/(1+\hat{d_{k}^{ui}})$ measures the confident performance similarity of property $g_{k}$ between graphs $G^{u}$ and $G^{i}$.
Adaptive Knowledge Elicitation and Alignment. While $\bar{I}(g_{k})$ is grounded in the observations from benchmarks, their adaptation to unseen tasks requires finding the adaptive importance of property $w^{u}_{k}$, which cannot be observed without testing on $G^{u}$. Notably, such adaptive importance of property for a given graph commonly motivates specific GNN designs in the graph learning literature [54, 47], e.g., GraphSAGE [18] handles high-degree graphs. Therefore, we novelly propose to leverage LLMs—pre-trained on a vast corpus of graph learning literature—as meta-controllers to adaptively elicit $w^{u}_{k}$ for $G^{u}$ by aligning the task descriptions, filtered graph properties, and high-performing architecture patterns in textual space, followed by contextual reasoning [31]. This process ensures that the data-specific weights $w^{u}_{k}$ reflect the interpretable literature insights between graph properties, GNNs, and performance, aligning with general empirical evidence to construct trustworthy meta-knowledge that empowers LLMs to quickly design effective, data-aware models for unseen tasks.
Knowledge Pool Construction. Using the computed task similarity set $\mathcal{S}^{u}$ of $G^{u}$, we rank the benchmark graphs (Appendix A.3) and select the Top-$N_{s}$ most similar sources. The knowledge pool $\mathcal{K}$ for $G^{u}$ is then constructed by aggregating the top-performing models from these selected sources:
$\mathcal{K}=\bigcup_{G^{i}\in\text{Top-}{N_{s}}(\mathcal{S}^{u})}\{(G^{i},\{\theta_{im}^{*}\}_{m=1}^{N_{m}})\},$ (6)
The top-performing models $\{\theta_{i}^{*}\}$ act as a truncated prior $\mathbb{P}(\theta_{i}^{*})$, summarizing accumulated design knowledge from benchmark $G^{i}$, while task similarity $\mathcal{S}(G^{u},G^{i})$ serves as a gating module for solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$, indicating the effectiveness of $\theta_{i}^{*}$ to the unseen graph $G^{u}$.
3.3 Model Suggestion and Refinement
Overall, our formulation of Knowledge Pool $\mathcal{K}$ represents the approximated posterior $\mathbb{P}(\theta_{i}^{*}|\mathcal{F}(G^{u}))$ that prioritizes models that align with data-aware evidence. This Bayesian-inspired process ensures solid usage of constructed meta-knowledge for designing models in both one-shot and a few attempts. We present our detailed procedure in Algorithm 1 and prompt designs in code implementation.
Initial Model Suggestion. By supplying contextual meta-knowledge of high similarities to the unseen graph $G^{u}$ from $\mathcal{K}$, we leverage LLM as a surrogate meta-controller to samples $N_{s}$ initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ based on the approximated posterior: $\theta_{ui}\leftarrow\mathcal{LLM}_{IMS}(\mathcal{F}(G^{u},\bar{I}),\mathcal{K}_{i}),$ where $\mathcal{K}_{i}$ represents the specific knowledge from the $i$-th similar benchmark graph in $\mathcal{K}$. Compared to existing AutoGNNs, our method avoids any model training on the unseen before giving promising designs, combining fine-grained literature and empirical insights for immediate model delivery.
Knowledge-Driven Refinement. To further optimize initial model proposals, we design a structured, knowledge-driven refinement strategy inspired by expert-like exploration and exploitation tradeoffs, which is illustrated in Appendix A.4. Our local search strategy iteratively improves models using the constructed meta-knowledge of high similarities to the unseen from $\mathcal{K}$. This refinement process balances exploration (testing diverse configurations) and exploitation (focusing on high-potential candidates), leveraging prior knowledge in textual space to optimize model designs efficiently.
4 Experiments
4.1 Experimental Settings
Task and Datasets. We conduct experiments on the node classification using 8 benchmark datasets from NAS-Bench-Graph [35]: Cora, Citeseer, PubMed [37], CS, Physics, Photo, Computer [38], and ogbn-arXiv [20]. Besides, we include 3 extra datasets for out-of-distribution (e.g., heterophily) generalization: DBLP [3], Flickr [56], and Actor [34]. The statistics are summarized in Appendix B.1.
Baselines. We compare DesiGNN with 4 categories of 17 baselines: (1) Manually Designed GNNs: 7 GNNs [24, 42, 18, 52, 10, 2, 32] showcasing human experts’ efforts in designing tailored models for specific types of graphs; (2) Classical Automated Approaches: 3 typical automated search strategies [27, 36, 65], plus 2 advanced AutoGNN systems [15, 62], all operating within the same search space as DesiGNN for fairness; (3) Meta-based Approaches: 3 meta-based, semi-supervised methods [54, 35, 5] that recommend models from the most similar benchmark graphs or based on meta-knowledge after testing anchor models on the unseen task; (4) LLM-Based Approaches: 2 recent SOTA LLM-Based methods [43, 12] that leverage LLMs for iterative architecture search.
Evaluation Metrics. To ensure reliability, we report the average accuracy and standard deviation over 10 runs. For automated methods, we also analyze the best-so-far accuracy after validating 1-30 model proposals to measure short-run efficiency. To assess efficiency in the case of varying LLM API traffic, we set the basic time unit as the time to validate a model proposal on the unseen dataset.
Implementation Details. The DesiGNN code is implemented in PyTorch and LangChain (GPT-4). For clarity, Desi-Init refers to initial model proposals generated in Section 3.3, while DesiGNN denotes the full pipeline, including model refinement in Section 3.3. Our experiments use NAS-Bench-Graph [35] for accumulated knowledge and a unified search space. To prevent data leakage, when a benchmark dataset is treated as unseen, it is completely excluded from the knowledge retrieval process and anonymized, ensuring fair evaluation. Out-of-distribution datasets (DBLP, Flickr, Actor) have access to all benchmark graphs as knowledge sources. We have validated that LLMs, given only dataset descriptions, lack prior knowledge of effective architectures in the NAS-Bench-Graph search space in Appendix D.1. Hyperparameters and more details are in Appendix B.1.
4.2 Main Results
Table 2: Initial performance comparison (accuracy) of GNNs designed by different methods. $N_{s}=3$ for Desi-Init. Best overall results are in bold, and best among each category are underlined.
Type Model Cora Cit. Pub. CS Phy. Pho. Com. arX. DBL. Fli. Act.
Manual GNNs GCN 80.97(0.39) 69.90(1.26) 77.46(0.61) 88.65(0.57) 90.85(1.20) 89.44(0.48) 83.16(0.55) 71.08(0.16) 84.25(0.25) 54.27(0.14) 24.78(1.79)
GAT 80.83(0.47) 70.70(0.71) 75.93(0.26) 88.72(0.73) 89.47(1.14) 89.93(1.75) 81.35(1.26) 71.24(0.10) 84.98(0.15) 51.85(0.26) 26.91(1.09)
SAGE 79.47(0.31) 66.13(0.90) 75.50(1.14) 87.81(0.18) 91.43(0.29) 88.29(1.03) 81.46(0.73) 70.78(0.17) 85.41(0.06) 52.84(0.19) 30.13(0.70)
GIN 79.77(0.38) 63.30(1.26) 76.74(0.86) 81.08(3.09) 86.67(0.86) 87.37(1.01) 73.95(0.16) 61.33(0.70) 82.82(0.82) 51.08(0.08) 27.15(0.38)
ChebNet 79.40(0.57) 67.03(1.02) 75.13(0.49) 89.50(0.36) 89.75(0.87) 86.65(0.77) 79.10(2.26) 70.87(0.10) 84.84(0.21) 53.75(0.16) 30.46(0.77)
ARMA 78.33(0.69) 66.20(0.75) 75.00(0.51) 89.87(0.35) 88.88(1.09) 86.55(3.35) 78.47(0.57) 70.87(0.17) 84.31(0.30) 54.23(0.04) 31.29(0.43)
k-GNN 78.06(0.47) 30.97(3.56) 75.38(0.97) 83.81(0.58) 88.98(0.54) 86.45(0.21) 76.31(1.34) 63.18(0.38) 83.59(0.07) 51.18(0.33) 32.74(0.68)
Auto. NNI Random 77.87(2.41) 66.64(1.32) 74.16(1.68) 81.78(9.41) 90.59(0.94) 89.04(2.55) 76.61(3.56) 68.93(1.82) 76.45(7.53) 52.50(0.72) 30.90(3.75)
EA 78.23(1.04) 66.40(2.63) 72.88(2.11) 87.03(2.64) 88.07(2.41) 87.30(1.38) 77.56(6.42) 68.28(2.95) 85.13(0.38) 53.22(0.92) 31.69(2.95)
RL 73.44(8.11) 65.35(2.40) 75.44(1.24) 86.17(5.09) 88.15(4.24) 89.48(1.35) 77.70(3.07) 68.00(4.71) 84.18(0.50) 52.07(2.84) 31.72(4.60)
Auto. AutoGL GNAS 78.55(1.20) 63.25(5.87) 73.04(1.64) 86.04(7.88) 89.54(1.52) 87.27(2.96) 70.96(9.66) 69.94(1.71) 84.46(0.41) 54.67(0.54) 32.31(3.25)
Auto-GNN 78.58(2.18) 65.60(2.69) 76.07(0.77) 89.06(0.42) 89.26(1.51) 89.34(1.75) 77.49(3.41) 67.62(1.72) 84.67(0.62) 50.97(1.21) 30.18(4.67)
Meta -based Kendall 67.73 69.20 71.80 88.56 91.56 88.90 76.85 71.49 - - -
Overlap 79.36 67.30 71.80 88.56 89.95 90.37 76.85 71.68 - - -
AutoTrans. 80.24(0.72) 68.19(1.22) 76.04(0.86) 89.19(0.77) 91.19(1.04) 90.93(0.37) 81.77(1.25) 71.37(0.39) - - -
LLM -based GPT4GNAS 78.50(0.37) 67.46(0.76) 73.89(0.86) 89.26(0.38) 89.44(1.94) 89.12(2.26) 77.21(5.26) 68.98(1.22) 84.93(0.22) 52.47(0.10) 34.26(0.47)
GHGNAS 79.13(0.45) 67.35(0.44) 74.90(0.57) 89.15(0.81) 88.94(2.57) 89.42(1.99) 77.04(3.96) 69.66(1.28) 85.06(0.15) 52.48(0.23) 33.72(2.72)
Ours Desi-Init 80.31(0.00) 69.20(0.16) 76.60(0.00) 89.64(0.08) 92.10(0.00) 91.19(0.00) 82.20(0.00) 71.50(0.00) 85.56(0.24) 55.16(0.11) 34.41(0.48)
DesiGNN 81.77(0.40) 71.00(0.09) 77.57(0.29) 90.51(0.42) 92.61(0.00) 92.38(0.06) 84.08(0.66) 72.02(0.18) 85.89(0.21) 55.44(0.06) 37.57(0.62)
Initial Model Suggestions. First, we use Desi-Init in Table 2 to refer to the initial model suggestion part in Section 3.3, i.e., the first GNN suggested by LLM within seconds. While human-designed GNNs like GCN and GAT excel on popular datasets (e.g., Cora, Citeseer), they fail to generalize to less common heterophily datasets such as Flickr and Actor. In contrast, Desi-Init surpasses these overly specialized models, demonstrating robustness across diverse graph datasets. Notably, Desi-Init outperforms semi-supervised meta-based baselines (Kendall, Overlap, AutoTransfer), validating our contributions of establishing a proficient “graph-GNN-performance” meta-knowledge through unsupervised task similarity assessment and graph understanding with adaptive properties. Furthermore, compared to automated baselines, Desi-Init achieves superior performance across all 11 datasets, underscoring the strength of our knowledge-driven approach over conventional automated methods (lack domain-specific insights) and LLM-based methods (rely only on commonplace knowledge). Besides, DesiGNN’s initial model proposals rank in the top 5.77% of all possible architectures, as shown in Table 7 at Appendix B.2. These results confirm that Desi-Init provides high-quality, tailored model proposals with negligible computational cost by capturing and applying domain-specific knowledge effectively, i.e., alleviating the expertise and efficiency challenges.
Table 3: Computational resources (in #proposal validations) needed by baselines to reach DesiGNN’s 10-validation performance. means failing to achieve within 100 validations in at least 1 of 10 trials.
#Model Proposal Validations
Model Cora Cit. Pub. CS Phy. Pho. Com. arX. DBL. Fli.
GNAS 40.3 48.3 21.5 38.8 $\infty$ 66 34.8 55 69.4 77.2
Auto-. 21.8 $\infty$ 11.6 18.4 43 29.3 17 59.3 76.2 $\infty$
Rand. 21.7 $\infty$ 45 18.8 89 33 22 87 64.7 $\infty$
EA 24.7 $\infty$ 48.8 40.8 $\infty$ 23 51.3 72.8 $\infty$ 46.3
RL 82 $\infty$ 11.8 14.8 $\infty$ 84.5 48.2 37 38.8 99
G4G. 30-40 $\infty$ 40-50 10 $\infty$ 30-40 10-20 40-50 40-50 $\infty$
GHG. 20-30 $\infty$ 40-50 10 $\infty$ 20-30 20-30 30-40 $\infty$ $\infty$
Model Refinement and Short-run Efficiency. Building on the initial proposals, DesiGNN (full pipeline) can rapidly refine models to outperform all manually designed and automated GNNs across 11 datasets, demonstrating its efficient use of meta-knowledge to exploit the high potential range. Figure 4 highlights the short-run efficiency of DesiGNN, which achieves significant performance improvements with fewer model proposal validations compared to automated baselines (full results are presented in Table 8, 9, 10 and Figure 8, 9 at Appendix B.2). Unlike those methods requiring extensive validations to identify feasible models, DesiGNN emulates human-like refinement strategies, reducing computational costs and accelerating optimization. Specifically, Table 3 shows that other automated methods require significantly more model validations to match DesiGNN’s performance after just 10 validations. These results confirm that DesiGNN effectively addresses computational inefficiencies by leveraging accumulated knowledge to enable immediate feedback and rapid optimization, delivering consistently superior performance with minimal computational costs or under limited budgets.
Figure 4: Short-run performance of DesiGNN compared to stronger automated baselines.
Figure 5: Task similarities computed by empirical performance (target), two meta-based baselines, and ours. Kendall’s $\tau$ correlation coefficient with the empirical ground truth is shown in parentheses.
4.3 Effective Knowledge via Task Similarity
To assess whether our unsupervised task similarity in Equation 5 reflects the probability of successful knowledge transfer defined in Definition 1, we quantify the hit rate, defined as the tighter probability $\delta^{*}$ that the empirically most relevant benchmarks (i.e., those maximizing ($\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$) are included in the Top-$N_{s}$ benchmarks retrieved by $\mathcal{S}$ ($N_{s}$ is the tolerance). A higher hit rate directly
Table 4: Hit rate $\delta^{*}$ across unseen datasets.
Top-s Benchs $N_{s}$=1 $N_{s}$=2 $N_{s}$=3
Kendall 0 0.25 0.625
Overlap 0 0.375 0.625
Property Similar. (w/o LLMs) 0.375 0.5 0.625
Description Only (with LLMs) 0.262 0.5 0.575
Property Only (with LLMs) 0.375 0.6 0.725
Both (with LLMs) 0.275 0.5 0.6
supports the probabilistic claim in Definition 1, where $\mathcal{S}$ correlates with the likelihood of transferring high-performing models from $G^{i}$ to $G^{u}$. As shown in Table 4, our unsupervised task similarity based on graph properties achieves the highest hit rate across all tolerance settings, outperforming methods relying on Kendall or Top 5% similarities and also the purely statistical property similarity (without LLMs to conduct adaptive knowledge elicitation and alignment). This demonstrates that our performance-oriented $\mathcal{S}$, derived from key graph properties and adaptive weighting by LLMs, effectively captures the meta-knowledge required for data-aware model design. Besides, we observed that there is a negative effect when incorporating descriptive inputs in similarity assessment, suggesting that some semantic properties (e.g., being a citation graph) are less relevant in designing model architecture (case study in Figure 1(b) and Appendix D.2). We further visualize the pairwise similarities between graphs in Figure 3(b)5, showing that our unsupervised task similarity closely aligns with empirical performance similarity (the highest Kendall’s $\tau$ correlation). This further validates the claim that $\mathcal{S}$ approximates a pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$ well, reinforcing its role in supporting the Bayesian-inspired framework in Section 3.2.
More Results. Due to limited space, the ablation study on main experiments, graph understanding, model suggestion, and model refinement are presented in Appendix C.1, C.2, C.3, C.4, respectively. Various LLMs’ innate knowledge, reasoning capabilities, and prompts are studied in Appendix D.
5 Conclusion
We present DesiGNN, a knowledge-centered framework that demonstrates deep expertise for proficiently designing data-aware GNNs and eliminates the external noise by systematically converting accumulated model design experiences into fine-grained meta-knowledge well fitted to meta-learning with LLMs. Leveraging a performance-oriented task similarity grounded in empirically validated graph properties and adaptive knowledge elicitation via LLMs, DesiGNN bridges the gap between literature insights and practical, data-specific model deployment. DesiGNN integrates graph understanding, effective knowledge retrieval, and strategic model suggestion and refinement following a Bayesian perspective, enabling tailored GNN designs for diverse graphs in one shot or a few attempts. Extensive experiments demonstrate that DesiGNN expedites the design process by not only delivering top-tier GNN designs within seconds but also achieving superior performance with minimal costs.
Societal Impact
Our work contributes to the growing movement toward high-level automation in artificial intelligence by enabling the proficient, data-aware design of Graph Neural Networks (GNNs) through structured knowledge reuse and language model alignment. As AI systems become more complex and specialized, the ability to automate model design across unseen tasks without human expertise offers transformative potential for a wide range of real-world applications. DesiGNN lowers the barrier to entry for applying advanced graph-based machine learning, making it more accessible to scientists, engineers, and practitioners in domains such as biology, social networks, and recommendation systems. By replacing ad hoc manual tuning with systematic and interpretable design strategies, our framework may promote reproducibility, consistency, and scalability in the deployment of machine learning systems. More broadly, this work aligns with the vision of LLM-powered agents acting as autonomous scientific assistants—able to rapidly translate raw data into effective models—thereby accelerating innovation in scientific discovery, infrastructure optimization, and decision-making. We hope our contribution supports the responsible expansion of AI automation by enabling efficient and expert-level design processes at scale.
However, such high-level automation also warrants caution. While DesiGNN significantly reduces computational cost and expertise requirements, its efficacy may be affected by the quality and representativeness of the knowledge source, along with the reliability of LLMs’ inherent contextual reasoning capability. Additionally, while our method operates a level removed from direct applications, any downstream deployment of GNNs—e.g., in social network analysis or recommender systems—can have far-reaching societal implications such as privacy leakage or algorithmic bias if not handled responsibly.
Limitations and Future Work
While the current implementation focuses on node classification and leverages a fixed search space from open benchmarks, future work can develop more nuanced mappings between more diverse datasets, more delicate configuration spaces, and the corresponding performance on other tasks, e.g., link prediction. Notably, it is worthwhile to study how to select or synthesize benchmark datasets that are representative of their top-performing model patterns so that the knowledge base can maintain maximum knowledge efficacy and better mitigate the out-of-distribution challenges. Besides, our research focuses on enhancing the short-run effectiveness of automated GNN designs, leaving long-run knowledge-driven optimization with a more nuanced search tradeoff as an appealing direction. Another potential is to extend the currently finite search space to an unlimited design space. Looking ahead, there are also promising avenues for extending our knowledge-driven logic to broader machine-learning contexts beyond graph data, e.g., image and tabular data.
Appendix A Methodology
In this section, we provide a detailed explanation of the methodology behind our DesiGNN framework, including Graph Understanding (Section 3.1), Knowledge Retrieval (Section 3.2), and GNN Model Suggestion and Refinement Modules (Section 3.3).
A.1 Algorithm
The DesiGNN methodology is detailed in Algorithm 1. The algorithm is divided into three main phases: Graph Understanding, Knowledge Retrieval, and GNN Model Suggestion and Refinement. It also includes details on the Self-evaluation Mechanism, Empirical Filtering Mechanism, Adaptive Filtering Mechanism, Re-rank Mechanism, Controlled Exploration, Model Promotion Mechanism, Directional Exploitation, and Optimization.
(a) Property confidences across graphs.
(b) Average confidences of graph properties.
Figure 6: Empirical studies of graph property confidence as a stronger indicator of effective meta-knowledge in GNNs design.
Algorithm 1 DesiGNN Methodology
Input: Unseen Graph Dataset $G^{u}$
Data: Benchmark Datasets $\{G^{i}\}_{i=1}^{n}$, Benchmark Knowledge Base $\mathcal{H}_{G}:\Theta\to P$, Self-evaluation Bank $BE$, Task-aware LLMs $\mathcal{LLM}_{GDC}$, $\mathcal{LLM}_{IMS}$, $\mathcal{LLM}_{KDR}$, Empirical Filter $\mathcal{F}(\cdot,\cdot)$, #Properties $N_{f}$, Knowledge Pool size $N_{s}$, #Top models $N_{m}$, #Candidates $N_{c}$, stop criteria $maxTrials$
Output: Optimized GNN architecture $\theta^{*}_{u}$ and model parameters $\omega^{*}$
1:  Phase 1: Graph Understanding Module
2:  if $\mathcal{F}(\cdot,\cdot)$ needs initialization or update then
3:     Self-evaluation Mechanism
4:     for each anchor dataset $G^{i}$ and property $g_{k}\in BE$ do
5:        for each other dataset $G^{j}\in BE,G^{j}\neq G^{i}$ do
6:           Compute $d_{k}^{ij}=|g_{k}^{i}-g_{k}^{j}|$ and retrieve $p^{ij}\in BE$
7:        end for
8:        $\mathcal{SR}(i,k)=\text{argsort}(\{d_{k}^{ij}\mid j\neq i,G^{j}\in\mathbf{G}\},\text{ASC})$
9:        $\mathcal{ER}(i)=\text{argsort}(\{p^{ij}\mid j\neq i,G^{j}\in\mathbf{G}\},\text{DESC})$
10:        $I(G^{i},g_{k}):=\text{KendallCorr}\big{(}\mathcal{SR}(i,k),\mathcal{ER}(i)\big{)}$
11:     end for
12:     $\bar{I}(g_{k})=\frac{1}{n}\sum_{i=1}^{n}I(G^{i},g_{k})$
13:     Initialize/Update $\mathcal{F}(\cdot,\cdot)=\text{Top}_{N_{f}}(\{\bar{I}(g_{k})\})$
14:  end if
15:  Generate unseen dataset’s description with $\mathcal{F}(G^{u},\bar{I})$ and benchmark datasets’ descriptions with $\mathcal{F}(G^{i},\bar{I})$
16:  Phase 2: Knowledge Retrieval Module
17:  Compute $\mathcal{S}^{u}=\{\frac{1}{N_{f}}\sum_{k=1}^{N_{f}}w^{u}_{k}\cdot\frac{\bar{I}(g_{k})}{1+\hat{d_{k}^{ui}}}\}_{i=1}^{n}$; $w^{u}_{k}$ is determined by $\mathcal{LLM}_{GDC}\big{(}\mathcal{F}(G^{u},\bar{I}),\mathcal{F}(\{G^{i}\}_{i=1}^{n},\bar{I})\big{)}$
18:  Retrieve $\mathcal{K}=\bigcup_{G^{i}\in\text{Top-}{N_{s}}(\mathcal{S}^{u})}\{(G^{i},\{\theta_{im}^{*}\}_{m=1}^{N_{m}})\}$
19:  Phase 3: Initial Model Suggestion
20:  for each $\mathcal{K}_{i}\in\mathcal{K}$ do
21:     $\theta_{ui}\leftarrow\mathcal{LLM}_{IMS}(\mathcal{F}(G^{u},\bar{I}),\mathcal{K}_{i})$
22:     $p^{ui}=\mathcal{H}(\theta_{ui},\omega;G^{u})$
23:  end for
24:  $BE\leftarrow BE\cup\{(G^{u},\{g_{k}^{u}\},\{p^{ui}\})\}$
25:  Phase 4: Model Proposal Refinement
26:  $R=\text{argsort}(\{p_{ui}\},\text{DESC})$
27:  $\mathcal{K}_{1:N_{s}}=\{(G^{R[j]},\{\theta_{R[j]m}^{*}\}_{m=1}^{N_{m}})\}_{j=1}^{N_{s}}$
28:  $\Theta_{u}^{T}\leftarrow(\theta_{u}^{R[1]},p_{u}^{R[1]})$; $\theta^{*}_{u}=\theta_{u}^{T}[1]$
29:  for $t=N_{s}$ to $maxTrials$ do
30:     $C^{t}=\{\theta_{ui}^{t^{\prime}}\mid\text{Crossover}(\theta^{*}_{u},\mathcal{K}_{2:N_{s}})\}_{i=1}^{N_{c}}$
31:     $\theta_{u}^{t^{\prime}}=\text{argsort}(\{\mathcal{H}_{\mathcal{K}_{1}}(\theta_{ui}^{t^{\prime}})\mid\theta_{ui}^{t^{\prime}}\in C^{t}\},\text{DESC})[1]$
32:     $\theta_{u}^{t}\leftarrow\mathcal{LLM}_{MPR}(\theta_{u}^{t^{\prime}},\Theta^{T},\mathcal{K}_{1})$
33:     $p_{u}^{t}=\mathcal{H}(\theta_{u}^{t},\omega;G^{u})$
34:     $\Theta_{u}^{T}\leftarrow(\theta_{u}^{t},p_{u}^{t})$
35:     if $p_{u}^{t}>p_{u}^{*}$ then
36:        $\theta^{*}_{u},\omega^{*}=\theta_{u}^{t},\omega$
37:        $p_{u}^{*}=p_{u}^{t}$
38:     end if
39:  end for
40:  return $\theta^{*}_{u},\omega^{*}$
A.2 Graph Understanding Module
The Graph Understanding Module (Section 3.1) is designed to automatically analyze the graph properties to enhance task comprehension beyond mere user input and establish an empirical foundation for our meta-knowledge construction and unsupervised task similarity. The module initially prioritizes the most influential properties based on their confidences averaged across all benchmark datasets, which is illustrated in Figure 6(b). The selected features are then used to generate textual descriptions of graph datasets, which are subsequently employed to perform adaptive elicitation, align knowledge, and finally identify similar benchmark datasets to retrieve data-aware model design knowledge.
A.2.1 Graph Properties Pool
We employ a comprehensive set of $16$ graph properties by analyzing the graph learning literature [54, 47]. Illustrated in Table 1, the corresponding properties name from $g_{1}$ to $g_{16}$ are listed below (and are also in the initialized order based on their confidences $\bar{I}(g_{k})$): the average clustering coefficient, average betweenness centrality, density, average degree centrality, average closeness centrality, average degree, edge count, graph diameter, average shortest path length, assortativity, average eigenvector centrality, feature dimensionality, node count, node feature diversity, connected components, and label homophily. Together, they could capture the topology of graphs, which commonly motivates the specific design of GNNs in research. The underlined properties are measured on sampled subgraphs to reduce computational overhead and ensure scalability.
A.2.2 Statistical and Empirical Similarities
To empirically study the missing meta-level understanding about graph-GNN-performance, we start with 16 graph properties $\{g_{k}\}_{k=1}^{16}$ from the literature to infer graph similarity from two aspects: 1) the statistical distance ranking based on graph properties $g_{k}$, and 2) the empirical performance ranking of transferred top-performing GNNs from other graphs $\{G^{j}\mid j\neq i\}$ to anchor graph $G^{i}$.
Specifically, given two graphs $G^{i}$ and $G^{j}\in\mathbf{G}$, we first compute the statistical distance $d_{k}^{ij}=|g_{k}^{i}-g_{k}^{j}|$ between $G^{i}$ and $G^{j}$ based on $g_{k}$, i.e., the reverse ranking of $\{d_{k}^{ij}\}_{k=1}^{16}$ can assess the sectional topological similarities between graphs $G^{i}$ and $G^{j}$ after normalization. Then, to quantify empirical transferability, we let $p^{ij}$ denote the performance of transferring the top-performing GNN patterns summarized by LLMs from another graph $G^{j}$ to $G^{i}$ (details in Section 3.3). Based on $d_{k}^{ij}$ and $p^{ij}$, we define two similarity rankings of other graphs $\{G^{j}\mid j\neq i,G^{j}\in\mathbf{G}\}$ to graph $G^{i}$:
$\mathcal{SR}(i,k)$ $=\operatorname{argsort}\left(\{d_{k}^{ij}\mid j\neq i,\,G^{j}\in\mathbf{G}\},\,\text{ASC}\right),$ (7)
$\mathcal{ER}(i)$ $=\operatorname{argsort}\left(\{p^{ij}\mid j\neq i,\,G^{j}\in\mathbf{G}\},\,\text{DESC}\right)$ (8)
where $\mathcal{SR}(i,k)$ represents the statistical similarity ranking based on property $g_{k}$, and $\mathcal{ER}(i)$ represents the empirical ground truth of knowledge transferability when meta-learning with LLMs.
A.3 Knowledge Retrieval Module
As the knowledge source of our Knowledge Retrieval Module (Section 3.2), NAS-Bench-Graph [35] is a comprehensive benchmark designed to facilitate unified, reproducible, and efficient evaluations of graph neural architecture search (GraphNAS) methods. This benchmark encapsulates a well-defined search space and a rigorous evaluation protocol, which encompasses the training and testing of 26,206 unique graph neural network (GNN) architectures across nine diverse node classification datasets.
A.3.1 Benchmark Datasets
The benchmark utilizes nine graph datasets of varied sizes and types, including citation networks (Cora, Citeseer, PubMed) [37], co-authorship graphs (Coauthor CS and Coauthor Physics) [38], co-purchase networks (Amazon Computers and Amazon Photo) [38], and large-scale graphs like ogbn-arXiv [20] from the Open Graph Benchmark. These datasets are employed with fixed semi-supervised [35] splits to ensure consistent evaluation settings.
A.3.2 Search Space Design
The search space in NAS-Bench-Graph is constructed as a directed acyclic graph (DAG), representing various GNN architectures. It includes macro architectural choices constrained to 9 distinct patterns: [0,0,0,0], [0,0,0,1], [0,0,1,1], [0,0,1,2], [0,0,1,3], [0,1,1,1], [0,1,1,2], [0,1,2,2], and [0,1,2,3]. The operations include seven prominent GNN layer types—GCN [24], GAT [42], GraphSAGE [18], GIN [52], ChebNet [10], ARMA [2], and k-GNN [32]—plus Identity and fully connected layers for residual connections and non-graph structure use, respectively.
Figure 7: Conceptual illustration of controlled exploration and directional exploitation in GNN model suggestion and refinement.
Table 5: Statistics of the datasets.
Dataset #Vtx #Edges Feat. Cls Metric
Cora 2,708 5,429 1,433 7 Acc.
Citeseer 3,327 4,732 3,703 6 Acc.
PubMed 19,717 44,338 500 3 Acc.
CS 18,333 81,894 6,805 15 Acc.
Physics 34,493 247,962 8,415 5 Acc.
Photo 7,487 119,043 745 8 Acc.
Computers 13,381 245,778 767 10 Acc.
arxiv 169,343 1,166,243 128 40 Acc.
proteins 132,534 39,561,252 8 112 ROC
DBLP 17,716 105,734 1,639 4 Acc.
Flickr 89,250 899,756 500 7 Acc.
Actor 7,600 30,019 932 5 Acc.
A.4 GNN Model Suggestion and Refinement
As introduced in Section 3.3 and illustrated in Figure 2, our model design refinement cycle continues iteratively, with each step methodically enhancing the model architecture based on accumulated knowledge and feedback. This process integrates user requirements, retrieved top-performing models from $\mathcal{K}$, and search space description (e.g., macro architectures and operations). The detailed procedure is as follows: (0) Re-Ranking: Initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ are re-ranked based on their validation performance, and the best-performing proposal $\theta_{u1}$ serves as the refinement starting point. (1) Controlled Exploration: Configurations from other models in $\mathcal{K}$ are crossovered with $\theta_{u1}$ to generate $N_{c}$ new candidates. (2) Model Promotion Mechanism: Candidates are ranked based on their retrieved performances on the benchmark dataset of $\mathcal{K}_{1}$, with the most promising candidate $\theta_{u}^{t^{\prime}}$ advanced for further refinement at iteration $t$. (3) Directional Exploitation: The LLM meta-controller $\mathcal{LLM}_{KDR}$ mutates $\theta_{u}^{t^{\prime}}$ using user requirements, task descriptions, search space details, and previous training logs. (4) Evaluation and Update: Each refined candidate $\theta_{u}^{t}$ is validated and added to the optimization trajectory $\Theta^{T}$. The best-performing model is updated as $\theta_{u}^{*}$.
As conceptually illustrated in Figure 7, the search space is navigated in a controlled manner through simulated crossovers between $\theta^{*}$ and configurations from $\mathcal{K}_{2:N_{s}}$, ensuring that exploration remains within high-potential range. Among these generated candidates, the one that demonstrates the highest potential based on its performance on the closest benchmark dataset is identified as the most promising for the unseen dataset. This candidate is then subjected to further exploitation by the LLM, guided by the most valuable directional insights derived from $\mathcal{K}_{1}$. Our empirical results support the short-run effectiveness of this local search strategy in quickly refining models. Our retrieval-then-verified promotion strategy not only ensures a reliable and efficient selection of new proposals by leveraging pre-existing, empirically validated knowledge but also requires only a single evaluation per optimization step. This sharply contrasts with existing traditional and LLM-based methods that necessitate accumulating knowledge from scratch and running multiple evaluations in parallel, thereby reducing computational overhead and accelerating the refinement process.
Appendix B Experiments
B.1 Experimental Settings
Table 6: Summary of hyperparameters used for each dataset
Dataset #Pre-process #Post-process Dimension Dropout Optimizer Learning Rate Weight Decay Epochs
Cora 0 1 256 0.7 SGD 0.1 0.0005 400
Citeseer 0 1 256 0.7 SGD 0.2 0.0005 400
PubMed 0 0 128 0.3 SGD 0.2 0.0005 500
CS 1 0 128 0.6 SGD 0.5 0.0005 400
Physics 1 1 256 0.4 SGD 0.01 0 200
Photo 1 0 128 0.7 Adam 0.0002 0.0005 500
Computers 1 1 64 0.1 Adam 0.005 0.0005 500
ogbn-arxiv 0 1 128 0.2 Adam 0.002 0 500
ogbn-proteins 1 1 256 0 Adam 0.01 0.0005 500
DBLP 1 1 256 0.5 SGD 0.1 0.0005 300
Flickr 1 1 128 0.5 Adam 0.001 0.0005 300
Actor 1 1 128 0.5 Adam 0.005 0.0005 400
Table 7: Performance Ranking of Initial Model Proposals.
Performance Rank (%)
Avg. Cora Cit. Pub. CS Phy. Pho. Com arX.
5.765 5.899 3.324 4.926 11.036 0.053 3.144 7.353 10.383
B.1.1 Task and Datasets.
We evaluate our DesiGNN framework on eleven diverse graph datasets, including eight out of nine benchmark datasets { Cora [37], Citeseer [37], PubMed [37], CS [38], Physics [38], Photo [38], Computer [38], ogbn-arXiv [20] } from NAS-Bench-Graph [35] and three additional datasets { DBLP [3], Flickr [56], Actor [34] }. All datasets are used to conduct node classification tasks, and their statistics are summarized in Table 7. The reason for excluding ogbn-proteins [20] in evaluation is that NAS-Bench-Graph only recorded partial model performance mapping due to explosion and out-of-memory errors, making it hard to conduct benchmarking comparisons.
B.1.2 Training Hyperparameters
The training hyperparameters used in different datasets are summarized in Table 6. For the nine benchmark datasets, we use the same hyperparameters as the NAS-Bench-Graph [35] benchmark to ensure fair comparisons. These hyperparameters were selected based on random searches on 30 anchor GNN architectures. The hyperparameters for the three additional datasets are recommended by LLMs following the same procedure from Graph Understanding (Section 3.1) to Initial Model Suggestion (Section 3.3), except the retrieved model architecture knowledge (Section 3.2) is replaced by the tuned hyperparameters of the corresponding benchmark datasets [35]. All the hyperparameters are fixed across all experiments for fairness. All experiments are conducted on a single NVIDIA RTX 3080 GPU and any GPU that could run a conventional GNN training can work. Notably, DesiGNN’s initial model recommendation can be delivered without GPU.
Table 8: Best model performance (accuracy) after 10 model validations on the benchmark datasets.
ACC (STD) %
Type Model Cora Citeseer PubMed CS Physics Photo Computer arXiv DBLP Flickr Actor
Auto. AutoGL GNAS 80.89(0.35) 67.45(1.40) 76.36(0.50) 88.94(0.93) 89.55(2.17) 87.76(3.37) 75.70(5.54) 69.84(1.46) 84.67(0.29) 54.69(1.02) 35.12(1.71)
Auto-GNN 80.63(1.13) 68.45(0.72) 76.35(0.66) 88.46(2.52) 90.94(0.48) 90.08(1.63) 78.39(3.76) 67.78(1.91) 84.84(0.36) 51.27(2.01) 31.32(5.31)
Auto. NNI Random 80.79(0.60) 65.53(4.95) 75.79(0.79) 88.95(1.41) 90.68(0.27) 90.76(1.00) 76.95(3.00) 71.27(0.52) 82.68(5.89) 52.93(1.21) 33.33(3.35)
EA 81.27(0.59) 67.29(1.37) 74.92(0.88) 88.08(1.61) 91.05(0.64) 89.63(1.26) 80.46(1.88) 70.40(0.50) 85.15(0.33) 54.38(0.87) 33.83(2.65)
RL 80.28(0.77) 67.39(1.06) 75.39(1.34) 88.47(1.53) 90.70(0.90) 88.84(0.39) 77.48(3.90) 69.12(1.78) 84.46(0.23) 53.34(0.94) 33.61(2.43)
LLM -based GPT4GNAS 80.33(0.39) 69.37(0.01) 76.40(0.20) 90.24(0.06) 91.06(0.61) 91.51(0.78) 82.88(1.25) 70.85(0.22) 85.45(0.16) 55.12(0.16) 36.14(0.54)
GHGNAS 80.51(0.45) 69.30(0.22) 76.49(0.32) 90.28(0.29) 90.94(0.28) 91.64(0.47) 82.40(1.21) 70.88(0.38) 85.42(0.16) 55.00(0.18) 36.56(0.49)
Our DesiGNN 81.69(0.54) 70.58(0.46) 77.17(0.29) 90.22(0.48) 92.61(0.00) 92.25(0.21) 83.15(0.61) 71.92(0.15) 85.84(0.27) 55.20(0.09) 36.58(2.63)
Table 9: Best model performance (accuracy) after 20 model validations on the benchmark datasets.
ACC (STD) %
Type Model Cora Citeseer PubMed CS Physics Photo Computer arXiv DBLP Flickr Actor
Auto. AutoGL GNAS 81.35(0.23) 69.49(0.65) 76.94(0.21) 90.12(0.27) 91.64(0.52) 91.47(0.19) 83.15(1.23) 71.49(0.43) 85.51(0.07) 55.22(0.21) 36.82(0.59)
Auto-GNN 81.07(0.65) 69.78(0.52) 77.21(0.58) 90.36(0.35) 91.92(0.65) 91.52(1.16) 82.73(1.17) 71.50(0.30) 85.65(0.18) 54.99(0.18) 37.10(0.91)
Auto. NNI Random 80.97(0.45) 69.57(0.23) 76.69(0.15) 90.20(0.27) 91.80(0.31) 91.74(0.46) 83.21(0.61) 71.54(0.23) 85.37(0.13) 55.01(0.34) 37.31(0.58)
EA 81.27(0.59) 67.82(0.93) 76.33(0.70) 89.53(1.03) 91.38(0.51) 91.67(0.49) 82.16(1.17) 71.48(0.27) 85.64(0.25) 55.01(0.23) 36.73(1.48)
RL 80.53(0.48) 69.69(0.51) 77.33(0.62) 90.30(0.44) 91.48(0.51) 91.75(0.31) 82.46(0.99) 71.22(0.24) 85.44(0.16) 55.01(0.13) 37.52(0.31)
LLM -based GPT4GNAS 80.73(0.15) 69.43(0.07) 76.88(0.41) 90.44(0.06) 91.84(0.52) 91.79(0.43) 83.54(0.53) 71.31(0.27) 85.57(0.14) 55.12(0.16) 36.66(0.73)
GHGNAS 80.66(0.42) 69.44(0.06) 76.93(0.22) 90.42(0.06) 91.59(0.56) 92.03(0.39) 82.93(0.63) 71.44(0.31) 85.43(0.15) 55.06(0.11) 37.01(0.57)
Our DesiGNN 81.69(0.52) 70.87(0.09) 77.27(0.41) 90.41(0.39) 92.61(0.00) 92.29(0.05) 83.43(0.77) 71.98(0.16) 85.85(0.29) 55.32(0.14) 37.57(0.62)
Table 10: Best model performance (accuracy) after 30 model validations on the benchmark datasets.
ACC (STD) %
Type Model Cora Citeseer PubMed CS Physics Photo Computer arXiv DBLP Flickr Actor
Auto. AutoGL GNAS 81.35(0.23) 69.67(0.47) 76.97(0.22) 90.18(0.31) 91.92(0.43) 91.61(0.23) 83.23(1.13) 71.54(0.36) 85.58(0.08) 55.23(0.20) 37.23(0.56)
Auto-GNN 81.46(0.41) 69.78(0.52) 77.37(0.51) 90.50(0.20) 91.92(0.65) 91.97(0.40) 83.06(0.91) 71.61(0.35) 85.66(0.17) 55.20(0.15) 37.50(0.83)
Auto. NNI Random 81.17(0.44) 69.63(0.16) 77.25(0.44) 90.39(0.14) 91.81(0.31) 92.04(0.32) 83.62(0.63) 71.55(0.22) 85.48(0.15) 55.17(0.12) 37.39(0.59)
EA 81.27(0.59) 68.05(0.70) 76.51(0.83) 89.70(0.80) 91.60(0.65) 91.78(0.60) 83.00(0.84) 71.62(0.26) 85.66(0.24) 55.09(0.21) 37.44(1.86)
RL 80.91(0.24) 69.88(0.24) 77.33(0.62) 90.47(0.19) 91.83(0.38) 91.75(0.31) 82.48(0.96) 71.47(0.08) 85.52(0.23) 55.05(0.13) 37.52(0.31)
LLM -based GPT4GNAS 81.31(0.24) 69.43(0.07) 76.90(0.41) 90.44(0.06) 92.12(0.21) 92.21(0.13) 83.96(0.83) 71.67(0.44) 85.57(0.14) 55.12(0.16) 36.70(0.70)
GHGNAS 81.39(0.07) 69.64(0.38) 76.93(0.22) 90.42(0.06) 92.06(0.17) 92.38(0.02) 83.28(0.43) 71.84(0.03) 85.47(0.11) 55.06(0.11) 37.01(0.57)
Our DesiGNN 81.77(0.40) 71.00(0.09) 77.57(0.29) 90.51(0.42) 92.61(0.00) 92.38(0.06) 84.08(0.66) 72.02(0.18) 85.89(0.21) 55.44(0.06) 37.57(0.62)
B.1.3 Module Hyperparameters
The hyperparameters used in the DesiGNN framework are summarized as follows: the best empirical number of confident graph properties $N_{f}=8$ (without descriptional inputs), the size of the knowledge pool $N_{s}=3$ to balance tolerance with computation cost (consider benchmark datasets with Top-$3$ similarities to the unseen dataset), the number of top models $N_{m}=30$ (only top-performing model design examples), the number of candidates $N_{c}=30$, and the maximum number of trials $maxTrials=30$ for fast deployment scenario. These hyperparameters are selected based on the ablation studies in Section C.2,  C.3, and  C.4, which are then fixed across all main experiments.
B.2 Main Results
B.2.1 Performance Ranking of Initial Model Proposals
In this section, Table 7 presents the performance ranking of initial model proposals within the model space defined by NAS-Bench-Graph [35]. The average performance ranking of initial model proposals across all benchmark datasets is Top-$5.77\%$, indicating that the initial model proposals suggested by our DesiGNN-Init without any prior training are highly competitive.
B.2.2 Model Refinement and Short-run Efficiency
In this section, we present the complete results of the short-run experiments in Table 8 (after 10 model validations), Table 9 (after 20 model validations), and Table 10 (after 30 model validations). The performance trajectories of all automated baselines are presented in Figure 8 and Figure 9.
Table 11: The main ablation study on each component in GNN Model Suggestion and Refinement. * is our complete setting.
ACC (STD) %
Method Cora Citeseer PubMed CS Physics Photo Computers arXiv
Initial Proposal Performance
Property Only* 80.31 (0.00) 69.20 (0.16) 76.60 (0.00) 89.64 (0.08) 92.10 (0.00) 91.19 (0.00) 82.20(0.00) 71.50 (0.00)
Descriptional Only 80.31 (0.00) 69.26 (0.00) 75.71 (0.00) 89.53 (0.00) 88.42 (0.00) 91.17 (0.13) 81.79 (1.21) 71.20 (0.34)
Both 80.31 (0.00) 69.26 (0.00) 75.71 (0.00) 89.55 (0.03) 91.45 (0.78) 91.13 (0.07) 82.20(0.00) 71.50 (0.00)
w/o Knowledge 79.30 (0.00) 55.29 (0.00) 71.56 (0.00) 81.94 (0.00) 91.45 (0.00) 86.61 (0.00) 69.32 (0.00) 70.68 (0.00)
Best Performance After 30 Validations
All* 81.77 (0.40) 71.00 (0.09) 77.57 (0.29) 90.51 (0.42) 92.61 (0.00) 92.38 (0.06) 84.08 (0.66) 72.02 (0.18)
w/o Re-rank 81.77 (0.40) 69.77 (0.31) 77.40 (0.24) 90.47 (0.00) 92.61 (0.00) 92.38 (0.06) 83.90 (0.14) 71.87 (0.12)
w/o Promotion 81.04 (0.42) 70.33 (0.00) 77.00 (0.30) 90.19 (0.47) 92.71 (0.18) 92.22 (0.16) 82.85 (0.01) 71.99 (0.01)
w/o Exploration 81.61 (0.10) 70.34 (0.64) 77.40 (0.14) 90.24 (0.33) 92.61 (0.00) 91.87 (0.18) 83.74 (0.23) 71.99 (0.08)
w/o Knowledge 80.90 (0.00) 69.93 (0.00) 76.72 (0.00) 90.24 (0.00) 91.82 (0.00) 92.21 (0.00) 83.63 (0.00) 71.71 (0.00)
Appendix C Ablation Studies
To delve deeper into the pivotal designs of DesiGNN, we conducted ablation and hyperparameter studies on three key modules, including Graph Understanding (Section 3.1), Knowledge Retrieval (Section 3.2), and GNN Model Suggestion and Refinement (Section 3.3).
C.1 Main Ablation Study
We present the main ablation study on the Knowledge Retrieval and the GNN Model Suggestion and Refinement module in Table 11, quantifying the impact of the retrieved knowledge and each mechanism in our framework. The results from the initial model suggestion stage underscore the importance of designing a proper Graph Understanding method, demonstrating that relying solely on filtered graph properties in dataset descriptions outperforms any use of descriptional inputs. In the short-run performance evaluation after 30 model validations, the Re-rank mechanism significantly improves the knowledge-driven model refinement process when the original order of the Knowledge Pool $\mathcal{K}$ is incorrect. Additionally, the Model Promotion mechanism, which simulates the strategy of human experts refining a model based on accumulated knowledge, plays a crucial role in enhancing the efficacy of model refinement. Lastly, the Directional Exploration mechanism is empirically beneficial, as it leverages the most relevant model design knowledge and the in-context learning ability of LLMs to further refine the best candidate model promoted.
Table 12: Average ranks of different combinations across datasets. A higher rank (with 1 being the highest) corresponds to better performance and lower variability in the initial model proposals.
Combined Rank
Top-$N_{s}$ Desc. g=0 g=2 g=4 g=6 g=8 g=10 g=16
$N_{s}$=1 True 2.250 2.312 2.250 2.125 2.437 2.625 2.312
False 2.562 2.125 2.562 1.812 2.187 2.250 2.125
$N_{s}$=2 True 2.500 2.187 2.625 1.937 1.687 2.312 2.750
False 2.625 2.375 2.812 2.000 1.500 2.187 2.250
$N_{s}$=3 True 2.500 1.937 2.875 1.875 1.937 2.187 2.000
False 2.125 1.687 2.000 1.750 1.500 1.3750 1.687
C.2 Ablation Studies on Graph Understanding
In this section, we study the combined effect of the number of selected properties $N_{f}$ and the usage of descriptional inputs in understanding and comparing graph datasets, under varying sizes $N_{s}$ of the Knowledge Pool, which influences the number of initial proposals suggested by LLMs (Section 3.3). The average rankings of different combinations across datasets in Table 12 illustrate that while the optimal $N_{f}$ varies with different $N_{s}$ values, integrating descriptional inputs often diminishes performance, particularly when $N_{f}>6$ and $N_{s}=3$. This finding supports our notion of “artificial hallucinations,” where descriptive inputs can hinder the understanding of tasks when sufficient graph properties are provided. From another aspect, this result also validates our idea of analyzing graph properties to design tailored GNN models, which came from insights documented in the literature.
Table 13: The distribution of the best initial model among the Top-3 benchmark.
Is Best Initial Model (%)
Scenario Best Second Third
Descriptional Only 30.0 50.0 20.0
$N_{f}=6$ 50.0 22.5 27.5
$N_{f}=8$ 52.5 37.5 10.0
$N_{f}=10$ 42.5 40.0 17.5
Table 14: Average performance rank of transferring the Top-$N_{m}$ model designs.
Average Rank
Top-$N_{m}$ $N_{m}$=1 $N_{m}$=10 $N_{m}$=20 $N_{m}$=30 $N_{m}$=40 $N_{m}$=50
2.7077 2.8750 2.8438 2.6875 2.9231 2.8769
C.3 Ablation Studies on GNN Model Suggestion
This study extends from Section 3.1 and quantifies the advantage of using the two $N_{s}=2$ or three $N_{s}=3$ most similar benchmarks instead of solely the most similar one $s=1$ (no tolerance). Table 13 shows the distribution of the best initial proposal stems from the first, second, or third most similar benchmark. For $N_{f}=8$ without descriptive inputs, approximately $53\%$ of trials identify the optimal initial proposal from the most similar benchmark, with only $10\%$ benefiting from the third. This data supports a balanced trade-off between performance and efficiency. Notably, this study also serves as the motivation for us to design the Re-rank Mechanism in Model Proposal Refinement (Appendix A.4).
Moreover, we explore the optimal number $N_{m}$ of top-performing examples from each benchmark when establishing the Knowledge Pool and assess the impact of using poor examples. Table 14 shows that $N_{m}=30$ yields the best average performance across all dataset combinations ($2160$ records), with further analysis indicating that poor examples do not significantly affect performance outcomes (T-statistic $=-0.1262$, P-value $=0.8998$). We thus adopt $N_{m}=30$ without bad examples for an appropriate token length of the LLM prompt.
C.4 Ablation Studies on GNN Model Refinement
We investigate the impact of using the $\mathcal{K}_{1}$ Knowledge Pool as context for LLMs and the Re-rank Mechanism that is based on the performance ranking of initial proposals. We also study the best number of candidates and the effect of adopting different $N_{s}$ numbers of knowledge bases to guide the controlled search space exploration. Table 15 displays T-statistics for various configurations of Knowledge Pool usage and Re-rank Mechanism. For the row of w/o candidate, the True/False columns stand for w/o re-rank, and vice versa. While statistical significance is not reached (p-value $>0.05$), certain trends emerge:
  • When $N_{s}=3$ knowledge bases are used for a broader perspective, employing a Re-rank Mechanism consistently adds value. This pattern is intuitive as the Re-rank Mechanism will become more influential as the number of knowledge bases increases.
  • Broader exploration with $N_{s}=3$ and a larger number of candidates (e.g., $30$) is likely effective, as we also have a directional exploitation mechanism to refine the diversity of potentially immature solutions.
Given the complexity of these patterns, we further assess the performance ranking of GNN Model Refinement hyperparameter combinations across benchmarks to determine the most effective settings. Table 16 suggests that the best hyperparameter combination is $N_{s}=3$ and $N_{m}=30$ with both Re-rank Mechanism and knowledge-driven refinement strategy enabled.
Table 15: T-statistic of the ablation on the usage of knowledge-driven Exploration and Re-rank Mechanism with different candidate number and $N_{s}$.
Candidate Number= 10 20 30
w/o Re-rank or w/o Exploration
Ablation $N_{s}$ True False True False True False
w/o Exploration 2 -0.280 0.582 0.512 1.977 0.557 0.566
3 0.863 0.072 0.055 -0.164 -0.054 -0.358
w/o Re-rank 2 0.647 1.383 -0.334 1.232 -2.092 -0.852
3 -0.116 -0.808 -0.593 -0.591 -0.751 -1.904
Table 16: Average performance rank of the GNN Model Refinement hyperparameter combinations across benchmark.
Candidate Number= 10 20 30
w/o Re-rank
$N_{s}$ w/o Exploration True False True False True False
2 True 7.33 12 7.83 9 13.5 9.17
False 7.16 14.5 11.7 17.33 16.17 11.83
3 True 11.7 11.17 11.83 9.5 15.33 7.33
False 12.33 10.17 11.17 10.17 12.83 7
Appendix D Case Studies
D.1 Case Studies: Lack of Prior Knowledge in LLMs
The first study verifies the inherent knowledge gap illustrated in Figure 1(a): lack of prior knowledge in LLMs about the top-performing models of benchmark datasets within the NAS-Bench-Graph [35] model space (Section A.3). The first two columns in Table 20 are cases where the descriptive inputs or graph properties are directly sent to LLMs for model suggestions. The results demonstrate that LLMs are only able to suggest commonplace layer connections and operations based on user-provided descriptional inputs (left panel) or key graph properties (middle panel). In fact, we have tested all the benchmark datasets in NAS-Bench-Graph [35] with descriptive inputs, and we found only two macro lists and three operation lists that LLM would recommend regardless of the datasets:
  • Architecture: [0, 1, 2, 3] and [0, 0, 1, 3].
  • Operations: [’gcn’, ’gat’, ’sage’, ’skip’], [’gcn’, ’gat’, ’sage’, ’gcn’], and [’gcn’, ’gat’, ’gin’, ’sage’],
After providing the tailored knowledge retrieved from our framework to LLMs (right panel), it can be clearly seen that the model it recommends is specialized and has obvious performance improvements.
D.2 Case Studies: Artificial Hallucination in LLMs
The second case study in Table 21 and Figure 1(b) examines the phenomenon of “artificial hallucination”, or the external noise issue, in LLMs when comparing the similarities between the unseen dataset and the benchmark datasets, as detailed in Section 3.2. Figure 5 (leftmost) illustrates that the empirically most similar benchmark datasets to PubMed are CS, Physics, and Citeseer, in descending order of similarity. When employing our Graph Understanding method in Section 3.1, which leverages confident graph properties (right panel), the top three most similar datasets are Citeseer, CS, and Physics (align well with empirical ground truth). However, the current practice that relies solely on descriptive inputs (left panel) identifies Cora, Citeseer, and ogbn-arxiv as the top three, which does not align with the empirical ground truth. This discrepancy arises because LLMs overly rely on the shared characteristic of being citation graphs, assuming that citation datasets like PubMed, Cora, Citeseer, and ogbn-arxiv should have similar model design preferences. This case study demonstrates that relying solely on external descriptive inputs is insufficient to capture the similarities between datasets, which can overwhelm other pertinent information, leading to inaccurate task similarity assessment and incorrect knowledge retrieval.
Table 17: Performance comparison of initial model suggestion with different LLMs as meta-controller.
Dataset Cora Cit. Pub. CS Phy. Pho. Com. arX.
Llama2 80.26 68.90 75.83 89.84 90.75 91.35 80.27 70.95
GPT-4 80.31 69.20 76.60 89.64 92.10 91.19 82.20 71.50
Table 18: Performance comparison of model refinement with different LLMs as meta-controller.
Dataset Cora Cit. Pub. CS Phy. Pho. Com. arX.
Llama2 80.96 70.38 77.40 90.19 91.47 92.09 83.69 71.63
GPT-4 81.77 71.00 77.57 90.51 92.61 92.38 84.08 72.02
Table 19: Performance comparison of different prompt designs on GPT4GNAS.
Dataset Cora Cit. Pub. CS Phy. Pho. Com. arX.
Original 81.31 69.43 76.90 90.44 92.12 92.21 83.96 71.67
Our Prompt 81.38 70.06 77.03 90.22 92.16 92.08 83.78 71.72
D.3 Case Studies: Different LLMs
To study the effect of different LLMs as the meta-controller, we adopt an open-source LLM, Llama2-13b, to replace the GPT-4 meta-controller we originally employed in our paper. Table 17 shows the comparative results (averaged over five runs) on the initial model proposal, while Table 18 shows the comparative results (averaged over five runs) after 30 rounds of model validations. We can see that replacing GPT-4 with Llama2 results in a slight degradation of the performance of recommended models in both settings. This degradation is more consistent after the refinement of the model design. This suggests that GPT-4 is superior at proposing a reasonable next step from the trajectory and extra knowledge (optimization tasks), consistent with the existing study [53]. Notably, because the initial model proposal phrase relies more on the quality of our elicited knowledge, Llama2 can also achieve close performance to GPT-4.
D.4 Case Studies: Prompt Design Impact
We studied whether fine-tuning in prompt engineering has a positive or negative effect on our work. Since we did not adopt any prompt fine-tuning approach, we added an alternative experiment to compare whether there is any consistent difference in the performance of LLM-based baseline GPT4GNAS [43] between using our prompt design and its original design. This includes adding topological graph descriptions and replacing model space descriptions and general instructions. No knowledge-related design from our method is being transferred because it is our key novelty. After prompt replacement, GHGNAS [12] became the same as GPT4GNAS [43], so we only reported one of them. Table 19 below shows the comparative results (averaged over five runs) after 30 rounds of refinements. We can see that after replacing the prompt design of the baseline with our design, the impact on the quality of the recommended models is inconsistent. This study justifies that our method does not improve performance by fine-tuning the prompt; instead, the main source of our performance improvement arises from our tailored model design knowledge.
Figure 8: Short-run performance trajectories of DesiGNN compared to all the automated baselines after validating 1-30 proposals.
Figure 9: Short-run performance trajectories of DesiGNN compared to all the automated baselines after validating 1-30 proposals.
Table 20: Example LLM Response on Recommending Model for PubMed dataset
Section Response (w/ Descriptional Only) Response (w/ Properties Only) Response (w/ Knowledge)
Decision For the unseen dataset: (Architecture: [0, 0, 1, 3], Operations: [’gcn’, ’gat’, ’sage’, ’gcn’]) For the unseen dataset: (Architecture: [0, 0, 1, 3], Operations: [’gcn’, ’gat’, ’sage’, ’gin’]) For the unseen dataset: (Architecture: [0, 0, 0, 1], Operations: [’gcn’, ’gat’, ’arma’, ’graph’])
Reason 1. Architecture Choice ([0, 0, 1, 3]): …… 2. Operations Choice ([’gcn’, ’gat’, ’sage’, ’gcn’]): - First Operation - GCN (’gcn’): The Graph Convolutional Network (GCN) is effective for aggregating neighborhood information, which is essential in a citation graph where the connectivity pattern (citations) plays a crucial role in defining the node (paper) features. - Second Operation - GAT (’gat’): The Graph Attention Network (GAT) introduces an attention mechanism, allowing the model to focus on the most relevant parts of the input graph structure. This is particularly useful in citation graphs for emphasizing influential papers. - Third Operation - GraphSAGE (’sage’): GraphSAGE is adept at sampling and aggregating features from a node’s neighborhood, which can help in efficiently handling the local neighborhood structures and diversity in node degrees typical in citation graphs. - Fourth Operation - GCN (’gcn’): Another layer of GCN helps to further refine the feature representation by aggregating higher-order neighborhood information, which can be crucial for capturing the broader context in the graph. This combination and sequence of operations are designed to leverage both the local and global structural information in the graph, which is critical for a citation graph where both direct and indirect citations (connections) can influence the classification of a paper into topics. The use of both GCN and GAT helps in effectively capturing different levels of granularity in neighbor interactions, while GraphSAGE allows for efficient computation by focusing on a sampled subset of neighbors. The architecture’s mix of parallel and sequential processing ensures a comprehensive feature transformation pathway, enhancing the model’s ability to learn complex patterns in the data. 1. Architecture Choice ([0, 0, 1, 3]): …… 2. Operations Choice ([’gcn’, ’gat’, ’sage’, ’gin’]): - ’gcn’ (Graph Convolutional Network layer): Effective for capturing neighborhood information through a simple yet powerful convolutional approach. Given the average degree and edge count, ’gcn’ can efficiently aggregate local neighborhood features without being too computationally expensive. - ’gat’ (Graph Attention Network layer): Utilizes attention mechanisms to weigh the importance of nodes’ neighbors. This is particularly useful in graphs where nodes vary significantly in their roles and influence, as indicated by the centrality measures. - ’sage’ (GraphSAGE): Samples and aggregates features from a node’s neighborhood, which can handle the graph’s density and scale effectively. This method is adaptive and can learn from different-sized neighborhoods, which is beneficial given the variability in node connectivity. - ’gin’ (Graph Isomorphism Network layer): Designed to capture more nuanced graph structural information, which can be crucial for understanding deeper patterns in the graph, especially given the non-trivial graph diameter and clustering coefficient. This combination and sequence of operations are designed to first broadly and then selectively refine the feature representation of the nodes, considering both local and global structural properties of the graph. This should provide a robust model capable of handling the complexities and variabilities in the graph metrics provided. 1. Architecture Choice: …… 2. Operation Selection: …… - ’arma’ (Auto-Regressive Moving Average): ARMA layers have been chosen in various top models, suggesting their utility in capturing more complex dependencies in graph data than standard convolutional layers. - ’graph’ (k-GNN): The inclusion of a k-GNN operation in the final step can extend the model’s capability to capture higher-order interactions, which might be crucial for the specific characteristics of the unseen dataset. 3. Combination of Operations: The combination of GCN, GAT, and ARMA before a final GRAPH operation provides a robust sequence of transformations. Starting with GCN allows for a broad capture of neighborhood information, followed by the selective attention mechanism of GAT and the complex filtering of ARMA, culminating in a k-GNN layer that can capture higher-order dependencies. This sequence is designed to progressively refine and enhance the feature representations. 4. General Observations from Top Models: The selected operations and architecture are among the most recurrent in the top-performing models from similar datasets. This suggests that they are likely to be effective in handling the types of graph-structured data encountered in these benchmarks, making them a prudent choice for the unseen dataset. By synthesizing these observations, the recommended model design aims to balance complexity and depth of feature extraction and transformation, which is crucial for effectively processing graph-structured data. This design is expected to perform well on the unseen dataset by leveraging proven elements from similar successful models.
Perf. 75.40 71.56 76.60
Table 21: Example LLM Response on Dataset Comparison for PubMed dataset
Benchmark w/ Descriptional Inputs w/ Confident Graph Properties
Cora Similarity Score: 0.8 Reason: Both the unseen dataset and Cora are citation graphs with papers as vertices and citations as links. Both use bag-of-words features and have ground-truth topics as labels. The main differences are in the number of features and classes, with Cora having more features and classes. Similarity Score: 0.1 Reason: The unseen dataset has a significantly higher graph diameter, average closeness centrality, and average betweenness centrality compared to Cora. Additionally, the general graph metrics such as edge count, average degree, and density are quite different, indicating a low similarity.
Citeseer Similarity Score: 0.75 Reason: Similar to Cora, Citeseer is also a citation graph with similar characteristics to the unseen dataset. The differences are more pronounced in the number of features and classes, with Citeseer having significantly more features. Similarity Score: 0.6 Reason: Citeseer has a closer match in terms of graph diameter and average closeness centrality with the unseen dataset. The average betweenness centrality and general graph metrics like edge count and average degree also show some similarities, suggesting a moderate level of task similarity.
CS Similarity Score: 0.2 Reason: CS is a co-authorship graph, differing significantly in the type of vertices (authors vs. papers) and the nature of features (keywords vs. bag-of-words). The task and evaluation metric are similar, but the structural differences lower the similarity score. Similarity Score: 0.3 Reason: CS shares a similar graph diameter with the unseen dataset but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a moderate to low similarity.
Physics Similarity Score: 0.2 Reason: Similar to CS, Physics is a co-authorship graph with authors as vertices and co-author relationships as links, which is structurally different from the unseen dataset’s citation graph nature. Similarity Score: 0.2 Reason: Physics has a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and other general graph metrics like edge count and average degree, suggesting a low similarity.
Photo Similarity Score: 0.3 Reason: Photo is a co-purchase graph, which is different in nature from a citation graph. However, both use bag-of-words features and have a similar number of features, slightly increasing the similarity score. Similarity Score: 0.1 Reason: Photo, while having a similar graph diameter, differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a low similarity.
Computers Similarity Score: 0.3 Reason: Similar to Photo, Computers is also a co-purchase graph. The similarity is slightly higher due to a closer number of features and classes compared to the unseen dataset. Similarity Score: 0.1 Reason: Computers has a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics like edge count and average degree, indicating a low similarity.
ogbn-arxiv Similarity Score: 0.7 Reason: ogbn-arxiv is a citation graph similar to the unseen dataset, but it uses word embeddings for features instead of bag-of-words, and has a significantly different number of classes. Similarity Score: 0.2 Reason: ogbn-arxiv shares a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, suggesting a low similarity.
ogbn-proteins Similarity Score: -0.5 Reason: ogbn-proteins is structurally and fundamentally different from the unseen dataset, being a protein association graph with a multi-label binary classification task and ROC-AUC as the evaluation metric. Similarity Score: 0.1 Reason: ogbn-proteins has a different graph diameter and significantly different average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a very low similarity.

### v3

@@
 Proficient Graph Neural Network Design by Accumulating Knowledge on Large Language Models
-Abstract
+Abstract.
 High-level automation is increasingly critical in AI, driven by rapid advances in large language models (LLMs) and AI agents.
@@
 Achieving proficiency in designing data-aware models—defined as the meta-level capability to systematically accumulate, interpret, and apply data-specific design knowledge—remains challenging for existing automated approaches, due to their inefficient construction and application of meta-knowledge.
-To achieve the meta-level proficiency, we propose DesiGNN, a knowledge-centered framework that systematically converts past model design experiences into structured, fine-grained knowledge priors well fitted to meta-learning with LLMs.
+To achieve meta-level proficiency, we propose DesiGNN, a knowledge-centered framework that systematically converts past model design experience into structured, fine-grained knowledge priors well-suited for meta-learning with LLMs.
 To account for the inherent variability and external noise, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs.
-By constructing a solid meta-knowledge between unseen graph understanding and known effective architecture patterns, DesiGNN can deliver top-5.77% initial model proposals for unseen datasets within seconds, and achieve consistently superior performance with minimal search costs against baselines.
-1 Introduction
-Automation has become a defining trend in artificial intelligence, reshaping how complex tasks are executed at scale without human intervention or the requirement of deep expertise.
-From code generation [7, 9, 60] and agentic workflow orchestration [57, 30, 64] to the multi-task generalization of foundation models [29], this paradigm shift is transforming how expertise is embedded and operationalized in AI systems.
-Among various frontiers, the automatic design of machine learning models stands as one of the most fundamental and impactful challenges in this automation landscape.
-In the past, automated machine learning (AutoML) [44, 45, 48, 62] has over-emphasized model search specific to independent data and tasks from scratch, which does not conform to the cross-domain and cross-task automation trends.
-The recent breakthroughs in large language models (LLMs) [11, 4, 41] and AI agents [25, 19, 63] with strong academic backgrounds and contextual reasoning have further fueled expectations—could they autonomously generate tailored models for unseen tasks, thereby democratizing machine learning?
-Inspired by LLMs’ general reasoning and generation capability, recent studies [40] have leveraged LLMs to replace human involvement beyond traditional abilities of AutoML, such as task understanding [43, 12] and problem formulation [49].
-Despite democratizing complex processes of designing models, realizing highly data-aware automated model design, especially for data-sensitive domains like graph learning, remains elusive for LLMs.
-Unlike tasks in vision or language domains where LLMs excel in proposing generic solutions that could dominate all tasks and data properties, designing Graph Neural Networks (GNNs) [24, 59, 52, 10, 18, 42] for specific graph properties and graph learning tasks requires a nuanced understanding of how data topology influences model architecture [54, 47].
-For example, the homophily nature behind GCN’s aggregation has catastrophic effects when propagating information over heterophily graphs, where two nodes with different labels are more likely to be linked.
-Unfortunately, current LLM-based methods often fully rely on LLMs’ inherent knowledge and reasoning ability, lacking fine-grained design knowledge about the relationship between graph-specific properties and effective GNN configurations, while also being sensitive to external input noise.
-As observed in Figure 1(a), the LLM agents adopted in recent methods [43, 12, 49] tend to suggest generally well-performing yet overly generic architectures, inherently falling short in tailoring designs to unseen graph structures.
-Worse still, as shown in Figure 1(b), they are vulnerable to “artificial hallucinations”—being easily misled by external, irrelevant user descriptions (e.g., being a citation network) or blindly enumerated data statistics that do not accurately reflect the data-aware challenges in model design, resulting in misleading knowledge association and ineffective designs.
+By constructing a solid meta-knowledge between unseen graph understanding and known effective architecture patterns, DesiGNN can deliver top-5.77% initial model proposals for unseen datasets within seconds and achieve consistently superior performance with minimal search cost compared to baselines.
+Neural Architecture Search; Graph Neural Networks; Large Language Models
+journalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USA.booktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAisbn: 979-8-4007-2292-9/2026/02doi: 10.1145/3773966.3777982ccs: Computing methodologies Search methodologiesccs: Computing methodologies Neural networksccs: Computing methodologies Knowledge representation and reasoning
+1. Introduction
 (a) Inherent Knowledge Gap: LLMs propose overly generic GNNs when provided with purely semantic and property descriptions, unable to induce data-specific configurations.
 (b) External Noise: LLMs misinterpret task similarity by weighing more on the citation networks category, causing “artificial hallucinations” and ineffective knowledge transfers.
-Figure 1: Case studies illustrating inherent and external limitations of existing LLM-based methods.
-Achieving proficiency in designing data-aware models thus demands far more than general LLM-based model design methods, which calls for a meta-level capability to systematically accumulate, interpret, and apply fine-grained, data-specific design knowledge.
+Figure 1. Case studies illustrating inherent and external limitations of existing LLM-based methods.
+Automation has become a defining trend in artificial intelligence, reshaping how complex tasks are executed at scale without human intervention or the requirement of deep expertise.
+From code generation (Cheng et al., 2023; Choi et al., 2023; Zhang et al., 2023b) and agentic workflow orchestration (Zhang et al., 2024; Liu et al., 2023; Zhuge et al., 2024; Xu et al., 2025; Ouyang et al., 2025) to the multi-task generalization of foundation models (Liu et al., 2025), this paradigm shift is transforming how expertise is embedded and operationalized in AI systems.
+Among various frontiers, the automated design of machine learning models stands as one of the most fundamental and impactful challenges in this automation landscape.
+In the past, automated machine learning (AutoML) (Wang et al., 2022a, b; Wei et al., 2021; Zhou et al., 2022) has over-emphasized model search specific to independent data and tasks from scratch, which does not conform to the cross-domain and cross-task automation trends.
+The recent breakthroughs in large language models (LLMs) (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023; Yue et al., 2025a) and AI agents (Krishnan, 2025; Zhu et al., 2025; Yue et al., 2025b) with strong academic backgrounds and contextual reasoning have further fueled expectations—could they autonomously generate tailored models for unseen tasks, thereby democratizing machine learning?
+Inspired by LLMs’ general reasoning and generation capabilities, recent studies (Tornede et al., 2023) have leveraged LLMs to replace human involvement beyond traditional abilities of AutoML, such as task understanding (Wang et al., 2023a; Dong et al., 2023) and problem formulation (Wei et al., 2023).
+Despite democratizing the complex process of model design, realizing highly data-aware automated model design, especially in data-sensitive domains like graph learning, remains elusive for LLMs.
+Unlike tasks in vision or language domains, where LLMs excel in proposing generic solutions that could dominate all tasks and data properties, designing Graph Neural Networks (GNNs) (Kipf and Welling, 2016; Zhang and Chen, 2018; Xu et al., 2018; Defferrard et al., 2016; Hamilton et al., 2017; Veličković et al., 2017) for specific graph properties and graph learning tasks requires a nuanced understanding of how data topology influences model architecture (You et al., 2020; Wang et al., 2023b).
+For example, the homophily nature behind GCN’s aggregation has catastrophic effects when propagating information over heterophily graphs, where two nodes with different labels are more likely to be linked.
+Unfortunately, current LLM-based methods often rely entirely on LLMs’ inherent knowledge and reasoning, lacking fine-grained design knowledge of the relationships between graph-specific properties and effective GNN configurations, and are also sensitive to external input noise.
+As observed in Figure 1(a), the LLM agents adopted in recent methods (Wang et al., 2023a; Dong et al., 2023; Wei et al., 2023) tend to suggest architectures that are generally well-performing yet overly generic, inherently falling short in tailoring designs to unseen graph structures.
+Worse still, as shown in Figure 1(b), they are vulnerable to “artificial hallucinations”—being easily misled by external, irrelevant user descriptions (e.g., being a citation network) or blindly enumerated data statistics that do not accurately reflect the data-aware challenges in model design, resulting in misleading knowledge association and ineffective designs.
+Achieving proficiency in designing data-aware models thus demands far more than general LLM-based model design methods, which call for a meta-level capability to systematically accumulate, interpret, and apply fine-grained, data-specific design knowledge.
 Humans, akin to training a new Ph.D.
 student, face substantial hurdles in accumulating this complex knowledge, often undergoing a lengthy learning process to design tailored models proficiently.
-Meta-based AutoML methods represent initial efforts toward this goal computationally, seeking transferable insights across tasks via learned task embeddings [54, 1, 26] or surrogate predictors [50, 51].
-These methods, however, often rely on simple, predefined meta-features that may fail to generalize robustly when facing complex data diversity, thus bending the current automation trend.
-Extending this concept to graphs, AutoTransfer [5] has made the first promising strides by creating task-model repositories and embedding-based knowledge transfers to improve search efficiency.
-Nonetheless, it remains limited by predefined task embeddings and the rigid aggregation of prior tasks into random search, insufficiently capturing the fine-grained, graph-aware meta-knowledge documented in the vast graph learning literature [54, 47].
-To overcome these limitations, we propose DesiGNN, a knowledge-centered and proficiency-oriented framework for the automated design of data-aware GNNs for unseen tasks.
-DesiGNN systematically converts past model design experiences into structured, fine-grained knowledge priors that are well fitted to meta-learning with LLMs, novelly bridging the long-standing gap between unseen graph understanding and known effective architecture patterns in graph application research.
-To address the observed external noise from descriptive inputs and the inherent variability in graph properties, DesiGNN constructs an empirically validated meta-level understanding about graph-GNN-performance from benchmarks and adaptively aligns key properties with graph-specific insights elicited from the graph learning literature, using LLMs as knowledge aligners to enhance the constructed meta-knowledge.
+Meta-based AutoML methods represent initial efforts toward this goal computationally, seeking transferable insights across tasks via learned task embeddings (You et al., 2020; Achille et al., 2019; Le et al., 2022) or surrogate predictors (Wen et al., 2020; White et al., 2021).
+These methods, however, often rely on simple, predefined meta-features that may fail to generalize robustly to complex data diversity, thus bending the current automation trend.
+Extending this concept to graphs, AutoTransfer (Cao et al., 2023) has made the first promising strides by creating task-model repositories and embedding-based knowledge transfers to improve search efficiency.
+Nonetheless, it remains limited by predefined task embeddings and the rigid aggregation of prior tasks into random search, insufficiently capturing the fine-grained, graph-aware meta-knowledge documented in the vast graph learning literature (You et al., 2020; Wang et al., 2023b).
+To address these limitations, we propose DesiGNN, a knowledge-centered and proficiency-oriented framework for the automated design of data-aware GNNs for unseen tasks.
+DesiGNN systematically converts past model design experiences into structured, fine-grained knowledge priors well suited to meta-learning with LLMs, thereby novelly bridging the long-standing gap between unseen graph understanding and known effective architecture patterns in graph application research.
+To address the observed external noise from descriptive inputs and the inherent variability in graph properties, DesiGNN constructs an empirically validated meta-level understanding of graph-GNN-performance from benchmarks and adaptively aligns key properties with graph-specific insights elicited from the graph learning literature, using LLMs as knowledge aligners to enhance the constructed meta-knowledge.
 This hybridization of empirical knowledge and LLM-guided reasoning empowers DesiGNN to deliver immediate, data-aware model proposals without any training, efficiently refine these architectures based on accumulated insights, and ultimately deliver superior performance with minimal computational overhead.
 Our contributions are as follows:
-  • We empirically identify two key limitations in current LLM-based GNN design approaches: the detrimental external noise from descriptive inputs and the inherent gap in fine-grained design knowledge.
+  • We empirically identify two key limitations in current LLM-based GNN design approaches: the external noise from descriptive inputs and the inherent gap in fine-grained design knowledge.
 Motivated by these insights, we propose a knowledge-centered framework that enhances meta-level proficiency in designing tailored, data-aware GNNs for unseen tasks by systematically accumulating, interpreting, and applying fine-grained meta-knowledge.
-  • To account for the inherent variability of graph properties to GNN performance, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs, constructing a solid meta-knowledge between graph topological properties and high-performing GNN architectures.
-This enables DesiGNN to proficiently deliver immediate, data-aware GNN recommendations without any direct training on the unseen datasets.
-  • Extensive experiments across 8 benchmarks and 3 out-of-distribution datasets demonstrate that DesiGNN delivers initial GNN proposals ranking in the top 5.77% of all possible architectures within seconds.
+  • To account for the inherent variability of graph properties on GNN performance, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs, building a solid meta-knowledge between graph topology and high-performing GNN architectures.
+This enables DesiGNN to proficiently deliver immediate, data-aware GNN recommendations without any training on the unseen data.
+  • Extensive experiments across 3 out-of-distribution datasets and 8 benchmarks demonstrate that DesiGNN delivers initial GNN proposals ranking in the top 5.77% of all possible architectures within seconds.
 Moreover, its knowledge-driven refinement strategy rapidly optimizes these initial suggestions, consistently outperforming automated methods in both performance and efficiency.
-2 Related Work
-2.1 Automated GNNs
-Automated GNNs (AutoGNNs) [44, 15, 62] aim to automatically find an optimal GNN architecture $\theta$ for an unseen graph $G^{u}$ using a controller $\pi$ parameterized by $\alpha$, where the architecture $\theta$ with network weight $\omega$ can achieve best performance $\mathcal{H}(\theta,\omega;G^{u})$ on $G^{u}$.
-In AutoGNNs, the search spaces are categorized into intra-layer [14] and inter-layer [21] designs, with strategies including reinforcement learning [14], evolutionary algorithms [39], and differentiable search [46].
-However, as introduced in Section 1, these approaches lack proficiency and demand substantial computational resources [33], as they optimize configurations $\theta$ without leveraging prior knowledge and require extensive sampling from $\pi_{\alpha}(\theta)$.
-Recently, meta-learning [13, 6] approaches have sought to overcome these limitations by transferring insights from past tasks through learned task embeddings [54, 5, 1] or surrogate predictors [50, 51, 28].
-While these methods mark an initial step toward knowledge-driven AutoGNNs, they rely on coarse, predefined meta-features, often incapable of capturing the generalizable relationships between diverse data properties and effective model configurations.
+2. Related Work
+Figure 2.
+The illustration of our knowledge-centered DesiGNN pipeline for designing GNNs.
+The orange module is Graph understanding, the green module is Knowledge Retrieval, and the purple module is Model Suggestion and Refinement.
+2.1. Automated GNNs
+Automated GNNs (AutoGNNs) (Wang et al., 2022a; Gao et al., 2021; Zhou et al., 2022) aim to automatically find an optimal GNN architecture $\theta$ for an unseen graph $G^{u}$ using a controller $\pi$ parameterized by $\alpha$, where the architecture $\theta$ with network weight $\omega$ can achieve best performance $\mathcal{H}(\theta,\omega;G^{u})$ on $G^{u}$.
+In AutoGNNs, the search spaces are categorized into intra-layer (Gao et al., 2019) and inter-layer (Huan et al., 2021) designs, with strategies including reinforcement learning (Gao et al., 2019), evolutionary algorithms (Shi et al., 2022), differentiable search (Wang et al., 2021), and search-based model fine-tuning (Zhili et al., 2024).
+Similar strategies exist for the knowledge graph (Di and Chen, 2023; Shimin et al., 2021; Di et al., 2021; Iyer et al., 2022).
+However, these approaches lack proficiency and demand substantial computational resources (Oloulade et al., 2021), as they optimize configurations $\theta$ without leveraging prior knowledge and require extensive sampling from $\pi_{\alpha}(\theta)$.
+Recently, meta-learning (Finn et al., 2017; Chen et al., 2021) methods have sought to overcome these limitations by transferring insights from past tasks through learned task embeddings (Cao et al., 2023; Achille et al., 2019) or surrogate predictors (Wen et al., 2020; White et al., 2021; Li et al., 2023).
+While these methods mark an initial step toward knowledge-driven AutoGNNs, they rely on coarse, predefined meta-features that are often incapable of capturing generalizable relationships between diverse data properties and effective model configurations.
 NAS Benchmark for Designing GNNs.
-NAS-Bench-Graph [35] offers a standardized benchmark for GNN architecture search, comprising nine datasets and 26,206 unique architectures with associated performance records.
-Similarly, there are many other benchmarks created by pioneers in different domains [8].
-While these benchmarks could provide valuable insights into model-performance relationships, existing methods underutilize their potential for knowledge-driven design due to the absence of a data dimension, which is the foundation of knowledge transfer.
-Therefore, only a few methods [54, 5] attempted to design feasible models for unseen tasks using such model-performance knowledge.
-Notably, they rely predominantly on semi-supervised or supervised techniques to fit the missing associations with the data, i.e., require direct training on the unseen data.
-2.2 LLMs for Designing GNNs and the Limitations
-Inspired by LLMs’ ability to solve optimization tasks [53], recent studies explore LLMs for automating model design [40, 61, 22, 60, 55, 58, 17], focusing on managing task formulation, architecture generation, and optimization.
-Auto2Graph [49] democratizes the usage of AutoGNNs by translating descriptive user inputs (e.g., Cora [37] is a citation network) into configuration specifications, while GPT4GNAS [43] and GHGNAS [12] iteratively refine designs through crafted prompts, focusing on different graph geneity.
+NAS-Bench-Graph (Qin et al., 2022) offers a standard benchmark for GNN architecture search, comprising nine datasets and 26,206 unique architectures with associated performance records.
+Similarly, many other benchmarks have been created by pioneers across different domains (Chitty-Venkata et al., 2023).
+While these benchmarks could provide valuable insights into model-performance relationships, existing works underutilize their potential for knowledge-driven design because they lack a data dimension, which is the foundation of knowledge transfer.
+Therefore, only a few methods (Liu et al., ; You et al., 2020; Cao et al., 2023) attempted to design feasible models for unseen tasks using such model-performance knowledge.
+Notably, they rely primarily on semi-supervised or supervised techniques to fill in missing data associations, i.e., require direct training on unseen data.
+2.2. LLMs for Designing GNNs and Limitations
+Inspired by LLMs’ ability to solve optimization tasks (Yang et al., 2023), recent studies explore their use to automate model design (Zheng et al., 2023; Jawahar et al., 2023; Zhang et al., 2023b; Yu et al., 2023; Zhang et al., 2023a; Guo et al., 2024), focusing on managing task formulation, architecture generation, and optimization.
+Auto2Graph (Wei et al., 2023) democratizes the use of AutoGNNs by translating descriptive user inputs (e.g., Cora (Sen et al., 2008) is a citation network) into configuration specifications.
+GPT4GNAS (Wang et al., 2023a) and GHGNAS (Dong et al., 2023) iteratively refine model designs through crafted prompts, focusing on different graph geneity.
 Despite democratizing the usage, they suffer from: (1) Over-reliance on external descriptive user inputs that result in poor alignment between graph properties and architecture designs; (2) Raw LLMs inherently lack deep expertise in fine-grained graph-GNN-performance relationships, limiting their ability to suggest data-aware models; (3) Refining model designs solely based on feedback from testing models, lacking actionable, knowledge-driven search strategies.
-3 DesiGNN: Proficient Graph Neural Network Design
-Figure 2: The illustration of our knowledge-centered DesiGNN pipeline for designing GNNs.
-As discussed in Section 1, current LLM-based methods suffer from the lack of proficiency in designing data-aware GNNs, where inherent knowledge gaps and external noise are two major barriers.
-To formalize these challenges, we adopt a Bayesian perspective to represent the automated GNNs design as solving the posterior belief over a candidate architecture $\theta_{i}$ when facing an unseen graph $G^{u}$:
-$\mathbb{P}(\theta_{i}|G^{u})=\frac{\mathbb{P}(G^{u}|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(G^{u})},$ (1)
+3. Proficient Graph Neural Network Design
+As discussed in Section 1, current LLM-based methods lack proficiency in designing data-aware GNNs, with inherent knowledge gaps and external noise as major barriers.
+To formalize these challenges, we adopt a Bayesian perspective to represent the automated GNNs design as solving the posterior belief over candidate architecture $\theta_{i}$ when facing an unseen graph $G^{u}$:
+(1) $\mathbb{P}(\theta_{i}|G^{u})=\frac{\mathbb{P}(G^{u}|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(G^{u})},$
 where $\mathbb{P}(\theta_{i})$ is derived from a knowledge pool $\mathcal{K}$ accumulated from prior tasks $\{G^{i}\}$.
@@
 This proficiency denotes the capability to systematically accumulate, interpret, and apply fine-grained, data-aware meta-knowledge relating graph properties to effective GNN architectures.
-From a Bayesian perspective, such capability can be conceptually represented as solving the posterior belief $\mathbb{P}(\theta_{i}|G^{u})$ with minimal or even without testing candidate models $\{\theta_{i}\}$ on $G^{u}$.
-To enhance such meta-level proficiency without directly testing the unseen data, DesiGNN explicitly integrates empirically validated graph-property knowledge with adaptive LLM alignment to construct a trustworthy meta-level understanding $\mathcal{M}$ between graphs $\mathbf{G}$, GNNs $\mathbf{\Theta}$, and performances $\mathbf{P}$: $\mathcal{M}:(\mathbf{G},\mathbf{\Theta})\rightsquigarrow\mathbf{P}$ based on crucial insights from graph literature [54, 47].
-Figure 2 illustrates our Bayesian-inspired pipeline: Graph Understanding, Knowledge Retrieval, and Model Suggestion and Refinement, each component systematically addressing external noise of descriptive data inputs and inherent meta-knowledge gaps, followed by how to strategically apply the constructed meta-knowledge in designing data-aware models.
-We present our pseudo-code in Algorithm 1 in Appendix A.1.
-3.1 Connect Graph Topology to GNN Performance
+In our formulation, such capability can be conceptually represented as solving for the posterior belief $\mathbb{P}(\theta_{i}|G^{u})$, with minimal or even no testing of candidate models $\{\theta_{i}\}$ on $G^{u}$.
+To enhance such meta-level proficiency without directly testing the unseen data, DesiGNN explicitly integrates empirically validated graph-property knowledge with adaptive LLM alignment to construct a trustworthy meta-level understanding $\mathcal{M}$ between graphs $\mathbf{G}$, GNNs $\mathbf{\Theta}$, and performances $\mathbf{P}$: $\mathcal{M}:(\mathbf{G},\mathbf{\Theta})\rightsquigarrow\mathbf{P}$ based on crucial insights from graph literature (You et al., 2020; Wang et al., 2023b).
+Figure 2 illustrates our Bayesian-inspired pipeline: Graph Understanding, Knowledge Retrieval, and Model Suggestion and Refinement, each component systematically addressing external noise in descriptive data inputs and inherent meta-knowledge gaps, and how to strategically apply the constructed meta-knowledge in designing data-aware models.
 (a) Property confidence varies across graphs.
-(b) Descriptive inputs induce invalid task associations that deviate from empirical truth. DesiGNN corr.=0.52.
-Figure 3: Empirical studies of two barriers: inherent property variability and external noise.
-Performance-oriented modeling of discrete graph data is essential to quickly capture data priors from massive benchmarks to construct trustworthy meta-knowledge, a key challenge for LLMs due to their inability to naturally interpret graph-structured data [16].
-As discussed in Section 2.2, existing LLM-based methods rely heavily on user-provided task descriptions, which fail to capture the key graph topology (e.g., heterophily) that frequently inspires specific GNN designs in the vast graph learning literature [54, 47].
-Our studies in Figure 1(b),  3(b) and Appendix D.2 demonstrate that semantic or purely descriptive inputs (e.g., is a citation graph) frequently mislead LLMs, causing “artificial hallucination” and ineffective solutions.
+(b) Descriptive inputs induce invalid task associations that deviate from empirical truth. DesiGNN exhibits a better correlation of 0.52.
+Figure 3. Empirical studies of two barriers: inherent property variability and external noise.
+3.1. Connect Graph Topology to GNN Design
+Performance-oriented modeling of discrete graph data is essential to quickly capture data priors from massive benchmarks to construct trustworthy meta-knowledge, a key challenge for LLMs due to their inability to naturally interpret graph-structured data (Guo et al., 2023).
+As discussed in Section 2.2, existing LLM-based methods rely heavily on user-provided task descriptions, which fail to capture key graph topologies (e.g., heterophily) that often motivate specific GNN designs in the vast graph learning literature (You et al., 2020; Wang et al., 2023b).
+Our studies in Figures 1(b),  3(b), and Section 4.3 demonstrate that semantic or purely descriptive inputs (e.g., a citation graph) frequently mislead LLMs, leading to “artificial hallucination” and ineffective solutions.
 Such descriptive inputs introduce significant external noise, undermining the reliability of generic LLM suggestions when connecting graphs to model performance.
 Thus, it is necessary to establish an empirical foundation that captures fine-grained graph-GNN-performance mappings, enabling the construction of trustworthy meta-knowledge for data-aware model design.
-Table 1: 16 graph topological properties
-Clustering Betweenness Density Degree centrality
-Closeness Degree Edge count Graph diameter
-Shortest path Assortativity Eigenvector Dimensionality
-Node count Diversity Connected Label homophily
-To empirically study the missing meta-level understanding about graph-GNN-performance, we start with 16 graph properties $\{g_{k}\}_{k=1}^{16}$ from the literature (see Table 1 and Appendix A.2) to infer graph similarity from two aspects: 1) the statistical distance ranking $\mathcal{SR}(i,k)$ based on graph properties $g_{k}$, and 2) the empirical performance ranking $\mathcal{ER}(i)$ of transferred top-performing GNNs from other graphs $\{G^{j}\mid j\neq i\}$ to anchor graph $G^{i}$.
-The concrete definition is presented in Appendix A.2.2.
-To assess how reliably specific graph properties predict successful GNN design knowledge transfers, we introduce the concept of graph property confidence for $g_{k}$ on $G^{i}$ through Kendall’s $\tau$ Coefficient [23]:
-$I(G^{i},g_{k}):=\text{KendallCorr}\big{(}\mathcal{SR}(i,k),\mathcal{ER}(i)\big{)},$ (2)
+To empirically study the missing meta-level understanding about graph-GNN-performance, we start with 16 graph properties $\{g_{k}\}_{k=1}^{16}$ from the literature: clustering, betweenness, density, degree centrality, closeness, degree, edge count, graph diameter, shortest path, assortativity, eigenvector, dimensionality, node count, diversity, connected component, and label homophily.
+We infer graph similarity from two aspects: 1) the statistical distance ranking $\mathcal{SR}(i,k)$ based on graph properties $g_{k}$, and 2) the empirical performance ranking $\mathcal{ER}(i)$ of transferred top-performing GNNs from other graphs $\{G^{j}\mid j\neq i\}$ to anchor graph $G^{i}$.
+To assess how reliably specific graph properties predict successful GNN design knowledge transfers, we introduce the concept of graph property confidence for $g_{k}$ on $G^{i}$ through Kendall’s $\tau$ Coefficient (Kendall, 1938):
+(2) $I(G^{i},g_{k}):=\text{KendallCorr}\big(\mathcal{SR}(i,k),\mathcal{ER}(i)\big),$
 where $I(G^{i},g_{k})$ quantifies the correlation between the topological distance measured by $g_{k}$ and the empirical performance differences in transferring different top-performing GNNs from $\{G^{j}\mid j\neq i\}$ to $G^{i}$.
-A higher $I(G^{i},g_{k})$ indicates that $g_{k}$ is a stronger indicator of effective model design transfers, thus playing a critical role in meta-learning with LLMs to design data-aware models for unseen tasks.
+A higher $I(G^{i},g_{k})$ indicates that $g_{k}$ is a stronger indicator of effective model design transfers, thus playing a key role in meta-learning with LLMs to design data-aware models for unseen tasks.
+Variability in Property Confidence.
+To validate the reliability of graph properties for GNN performance, we assess 16 confidence scores $\{I(G^{i},g_{k})\}_{k=1}^{16}$ for each anchor graph $G^{i}$ against the other 8 graphs in the benchmark (Qin et al., 2022).
+Figure 3(a) shows that confidence $I(G^{i},g_{k})$ fluctuates widely across diverse graphs.
+e.g., while $I(\text{Physics},g_{16}\text{=homophily})$ is positive, the same property exhibits negative correlations for other graphs.
+This inherent variability highlights that specific graph properties are strong indicators of transferability for certain graphs but unreliable for others.
+Notably, this directly exposes why generic LLM suggestions fail to capture fine-grained insights—LLMs cannot implicitly discern when particular properties matter without explicit, empirically validated priors.
+Empirical Filtering as Trustworthy Foundation.
+Recognizing those empirical insights on external noise and inherent property variability, we first use empirical filtering to preliminarily identify graph properties that are reliably influential by averaging their predictive confidences $I(G^{i},g_{k})$ across benchmark graphs as $\bar{I}(g_{k})=\frac{1}{n}\sum_{i=1}^{n}I(G^{i},g_{k}).$ This process ensures that our property set built from the literature is also empirically grounded, reducing susceptibility to external noise and inherent property variability.
+Then, based on the empirical study in Figure 4, we develop a filter $\mathcal{F}(G^{i},\bar{I})$ to correlate the discrete graph modeling with transferred model performance, selecting the Top-$N_{f}$ properties with the highest $\bar{I}(g_{k})$.
+This leads to the alternative formulation of Equation 1:
+(3) $\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))=\frac{\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(\mathcal{F}(G^{u}))}.$
+This mechanism mitigates the typical LLM failure of “artificial hallucinations”—often triggered by unprofessional graph modeling—at negligible offline computational cost.
+To accumulate new knowledge, we design a self-evaluation bank $BE$ that stores measured graph properties and similarity rankings for new graphs on the fly, enabling $\mathcal{F}$ to scalably evolve over time by integrating new empirical evidence.
+Grounded in empirical transferability, this mechanism serves as a trustworthy foundation for capturing graph-to-graph ($G^{i}$ to $G^{j}$) meta-priors that underpin the correlations between graphs and high-performing GNN models, yet it still calls for an effective property variability handler.
+Figure 4. Average confidence of graph properties.
+3.2. Knowledge Retrieval Module
+Proficiency in GNN design largely depends on capturing meta-knowledge in the form of the posterior belief for unseen tasks.
+However, even after empirically filtering and establishing a general correlation between graph properties and model performance, solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ remains a resource-intensive accumulation of testing various $\theta_{i}$ on $G^{u}$.
+To overcome it, we novelly propose to leverage the known likelihood $\mathbb{P}(\mathcal{F}(G^{i})|\theta_{i})$ on benchmark datasets to form a practical proxy for $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ on the unseen.
+Eliciting such fine-grained model design knowledge paves the way for an efficient approximation of the posterior $\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))$ with accumulated prior knowledge.
+Finding such a proxy for the unseen task among the massive benchmark knowledge requires an unsupervised task similarity formulation that properly assesses the likelihood of successful model design knowledge transfer from $G^{i}$ to $G^{u}$, which has to reflect the performance outcomes $\mathcal{H}(\cdot,\cdot;G^{u})$.
+Definition 3.1 (Task Similarity for Model Design Knowledge Transfer).
+Given an unseen graph $G^{u}$ and two benchmark graphs $G^{i}$ and $G^{j}$, each associated with $N_{m}$ top-performing models $\{\theta^{*}_{m}\}_{m=1}^{N_{m}}\in\Theta$, an effective task similarity $\mathcal{S}(\cdot,\cdot)$ for transferring model design knowledge satisfies:
+(4) $\mathcal{S}(G^{u},G^{i})\geq\mathcal{S}(G^{u},G^{j})$
+$\Rightarrow$ $\mathbb{P}\Bigl[\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})\geq\mathcal{H}(\{\theta^{*}_{jm}\},\cdot,G^{u})\Bigr]\geq\delta.$
+where $\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$ denotes the performance of the transferred top-performing pattern $\{\theta^{*}_{im}\}$ from benchmark $G^{i}$ on the unseen $G^{u}$, and $\delta$ is a confidence threshold.
+This ensures that higher task similarity between $G^{u}$ and $G^{i}$ implies a larger likelihood of high-performing transfer from $G^{i}$ to $G^{u}$.
+To design unsupervised task similarity between $G^{u}$ and $G^{i}$, we first apply the filter $\mathcal{F}(\cdot)$ on $G^{u}$ and $G^{i}$ to establish an empirical foundation based on performance-oriented confidence $\bar{I}(g_{k})$.
+However, as discussed in Section 3.1, graph properties inherently exhibit varying levels of confidence across graphs, making it challenging to transfer knowledge reliably with standalone statistical properties.
+To account for inherent variability, we further introduce the adaptive importance $w^{u}_{k}$ for each property $g_{k}$ specifically tailored to the unseen graph $G^{u}$ and define our unsupervised task similarity as:
+(5) $\mathcal{S}(G^{u},G^{i})=\frac{1}{N_{f}}\sum_{k=1}^{N_{f}}w^{u}_{k}\cdot\frac{\bar{I}(g_{k})}{1+\hat{d_{k}^{ui}}}$,
 […]
-Empirical Filtering as Trustworthy Foundation.
-Recognizing those empirical insights on external noise and inherent property variability, we first utilize empirical filtering to preliminarily identify reliably influential graph properties by averaging their predictive confidences $I(G^{i},g_{k})$ across benchmark graphs as $\bar{I}(g_{k})=\frac{1}{n}\sum_{i=1}^{n}I(G^{i},g_{k}).$ This process ensures that our property set built from the literature is also empirically grounded, reducing susceptibility to external noise and inherent property variability.
-Then, based on the empirical observation in Figure 6(b), we develop a filter $\mathcal{F}(G^{i},\bar{I})$ to correlate the modeling of discrete graph data with transferred model performance, selecting the Top-$N_{f}$ properties with the highest $\bar{I}(g_{k})$.
-This leads to the alternative formulation of Equation 1:
-$\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))=\frac{\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(\mathcal{F}(G^{u}))}.$ (3)
-This mechanism mitigates the typical LLM failure of “artificial hallucinations” often triggered by unprofessional graph modeling with negligible computational cost offline.
-To accumulate new knowledge, we design a self-evaluation bank $BE$, which stores measured graph properties and similarity rankings for new graphs on the fly, allowing $\mathcal{F}$ to scalably evolve over time by integrating new empirical evidence.
-Grounded in empirical transferability, this mechanism serves as a trustworthy foundation for capturing graph-to-graph ($G^{i}$ to $G^{j}$) meta-priors that underpin the correlations between graphs and high-performing GNN models, but still calling for an effective property variability handler.
-3.2 Knowledge Retrieval Module
-Proficiency in GNN design largely depends on capturing the meta-knowledge in the form of the posterior belief for unseen tasks.
-However, even after empirically filtering and establishing a general correlation between graph properties and model performance, solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ remains a resource-intensive accumulation of testing various $\theta_{i}$ on $G^{u}$.
-To overcome it, we novelly propose to utilize the known likelihood $\mathbb{P}(\mathcal{F}(G^{i})|\theta_{i})$ on the benchmark datasets to form a practical proxy for $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ on the unseen.
-Eliciting such fine-grained model design knowledge paves the way for efficient approximation of the posterior $\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))$ with accumulated prior knowledge.
-Finding such a practical proxy for the unseen task among the massive benchmark knowledge requires an unsupervised task similarity formulation that properly assesses the likelihood of successful model design knowledge transfer from $G^{i}$ to $G^{u}$, which has to reflect the performance outcomes $\mathcal{H}(\cdot,\cdot;G^{u})$.
-Definition 1 (Task Similarity for Model Design Knowledge Transfer)
-Given an unseen graph $G^{u}$ and two benchmark graphs $G^{i}$ and $G^{j}$, each associated with $N_{m}$ top-performing models $\{\theta^{*}_{m}\}_{m=1}^{N_{m}}\in\Theta$, an effective task similarity $\mathcal{S}(\cdot,\cdot)$ for transfering model design knowledge satisfies:
-$\mathcal{S}(G^{u},G^{i})\geq\mathcal{S}(G^{u},G^{j})$ (4)
-$\Rightarrow$ [($\mathbb{P}\Bigl{[}\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})\geq\mathcal{H}(\{\theta^{*}_{jm}\},\cdot,G^{u})\Bigr{]}\geq\delta.$
-where ($\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$ denotes the performance of the transferred top-performing pattern $\{\theta^{*}_{im}\}$ from benchmark $G^{i}$ on the unseen $G^{u}$, and $\delta$ is a confidence threshold.
-This ensures that higher task similarity between $G^{u}$ and $G^{i}$ implies a larger likelihood of high-performing transfer from $G^{i}$ to $G^{u}$.
-To design such performance-oriented task similarity between $G^{u}$ and $G^{i}$ unsupervisedly, we first apply the filter $\mathcal{F}(\cdot)$ (initialized with accumulated benchmark knowledge in Section 3.1) on $G^{u}$ and $G^{i}$ to establish an empirical foundation based on performance-oriented confidence $\bar{I}(g_{k})$.
-However, as discussed in Section 3.1, graph properties inherently exhibit varying levels of confidences across graphs, making it challenging to transfer knowledge reliably with standalone statistical properties.
-To account for this inherent variability, we further introduce the adaptive importance $w^{u}_{k}$ for each property $g_{k}$ specifically tailored to the unseen graph $G^{u}$ and formulate our unsupervised task similarity as:
-$\mathcal{S}(G^{u},G^{i})=\frac{1}{N_{f}}\sum_{k=1}^{N_{f}}w^{u}_{k}\cdot\frac{\bar{I}(g_{k})}{1+\hat{d_{k}^{ui}}},$ (5)
-where $N_{f}$ is the number of influential graph properties selected via empirical filtering and $\bar{I}(g_{k})/(1+\hat{d_{k}^{ui}})$ measures the confident performance similarity of property $g_{k}$ between graphs $G^{u}$ and $G^{i}$.
 Adaptive Knowledge Elicitation and Alignment.
 While $\bar{I}(g_{k})$ is grounded in the observations from benchmarks, their adaptation to unseen tasks requires finding the adaptive importance of property $w^{u}_{k}$, which cannot be observed without testing on $G^{u}$.
-Notably, such adaptive importance of property for a given graph commonly motivates specific GNN designs in the graph learning literature [54, 47], e.g., GraphSAGE [18] handles high-degree graphs.
-Therefore, we novelly propose to leverage LLMs—pre-trained on a vast corpus of graph learning literature—as meta-controllers to adaptively elicit $w^{u}_{k}$ for $G^{u}$ by aligning the task descriptions, filtered graph properties, and high-performing architecture patterns in textual space, followed by contextual reasoning [31].
-This process ensures that the data-specific weights $w^{u}_{k}$ reflect the interpretable literature insights between graph properties, GNNs, and performance, aligning with general empirical evidence to construct trustworthy meta-knowledge that empowers LLMs to quickly design effective, data-aware models for unseen tasks.
+Notably, such adaptive importance of property for a given graph commonly motivates specific GNN designs in the graph learning literature (You et al., 2020; Wang et al., 2023b), e.g., GraphSAGE (Hamilton et al., 2017) handles high-degree graphs.
+Therefore, we novelly propose to leverage LLMs—pre-trained on a vast corpus of graph learning literature—as meta-controllers to adaptively elicit $w^{u}_{k}$ for $G^{u}$ by aligning task descriptions, filtered graph properties, and high-performing architecture patterns in textual space, followed by contextual reasoning (Min et al., 2022).
+This process ensures that the data-specific weights $w^{u}_{k}$ reflect interpretable insights from the literature on graph properties, GNNs, and performance, aligning with general empirical evidence to construct trustworthy meta-knowledge that empowers LLMs to quickly design effective, data-aware models for unseen tasks.
 Knowledge Pool Construction.
-Using the computed task similarity set $\mathcal{S}^{u}$ of $G^{u}$, we rank the benchmark graphs (Appendix A.3) and select the Top-$N_{s}$ most similar sources.
-The knowledge pool $\mathcal{K}$ for $G^{u}$ is then constructed by aggregating the top-performing models from these selected sources:
-$\mathcal{K}=\bigcup_{G^{i}\in\text{Top-}{N_{s}}(\mathcal{S}^{u})}\{(G^{i},\{\theta_{im}^{*}\}_{m=1}^{N_{m}})\},$ (6)
-The top-performing models $\{\theta_{i}^{*}\}$ act as a truncated prior $\mathbb{P}(\theta_{i}^{*})$, summarizing accumulated design knowledge from benchmark $G^{i}$, while task similarity $\mathcal{S}(G^{u},G^{i})$ serves as a gating module for solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$, indicating the effectiveness of $\theta_{i}^{*}$ to the unseen graph $G^{u}$.
-3.3 Model Suggestion and Refinement
-Overall, our formulation of Knowledge Pool $\mathcal{K}$ represents the approximated posterior $\mathbb{P}(\theta_{i}^{*}|\mathcal{F}(G^{u}))$ that prioritizes models that align with data-aware evidence.
-This Bayesian-inspired process ensures solid usage of constructed meta-knowledge for designing models in both one-shot and a few attempts.
-We present our detailed procedure in Algorithm 1 and prompt designs in code implementation.
-Initial Model Suggestion.
-By supplying contextual meta-knowledge of high similarities to the unseen graph $G^{u}$ from $\mathcal{K}$, we leverage LLM as a surrogate meta-controller to samples $N_{s}$ initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ based on the approximated posterior: $\theta_{ui}\leftarrow\mathcal{LLM}_{IMS}(\mathcal{F}(G^{u},\bar{I}),\mathcal{K}_{i}),$ where $\mathcal{K}_{i}$ represents the specific knowledge from the $i$-th similar benchmark graph in $\mathcal{K}$.
-Compared to existing AutoGNNs, our method avoids any model training on the unseen before giving promising designs, combining fine-grained literature and empirical insights for immediate model delivery.
-Knowledge-Driven Refinement.
-To further optimize initial model proposals, we design a structured, knowledge-driven refinement strategy inspired by expert-like exploration and exploitation tradeoffs, which is illustrated in Appendix A.4.
-Our local search strategy iteratively improves models using the constructed meta-knowledge of high similarities to the unseen from $\mathcal{K}$.
-This refinement process balances exploration (testing diverse configurations) and exploitation (focusing on high-potential candidates), leveraging prior knowledge in textual space to optimize model designs efficiently.
-4 Experiments
-4.1 Experimental Settings
-Task and Datasets.
-We conduct experiments on the node classification using 8 benchmark datasets from NAS-Bench-Graph [35]: Cora, Citeseer, PubMed [37], CS, Physics, Photo, Computer [38], and ogbn-arXiv [20].
-Besides, we include 3 extra datasets for out-of-distribution (e.g., heterophily) generalization: DBLP [3], Flickr [56], and Actor [34].
-The statistics are summarized in Appendix B.1.
-Baselines.
-We compare DesiGNN with 4 categories of 17 baselines: (1) Manually Designed GNNs: 7 GNNs [24, 42, 18, 52, 10, 2, 32] showcasing human experts’ efforts in designing tailored models for specific types of graphs; (2) Classical Automated Approaches: 3 typical automated search strategies [27, 36, 65], plus 2 advanced AutoGNN systems [15, 62], all operating within the same search space as DesiGNN for fairness; (3) Meta-based Approaches: 3 meta-based, semi-supervised methods [54, 35, 5] that recommend models from the most similar benchmark graphs or based on meta-knowledge after testing anchor models on the unseen task; (4) LLM-Based Approaches: 2 recent SOTA LLM-Based methods [43, 12] that leverage LLMs for iterative architecture search.
-Evaluation Metrics.
-To ensure reliability, we report the average accuracy and standard deviation over 10 runs.
-For automated methods, we also analyze the best-so-far accuracy after validating 1-30 model proposals to measure short-run efficiency.
-To assess efficiency in the case of varying LLM API traffic, we set the basic time unit as the time to validate a model proposal on the unseen dataset.
-Implementation Details.
-The DesiGNN code is implemented in PyTorch and LangChain (GPT-4).
-For clarity, Desi-Init refers to initial model proposals generated in Section 3.3, while DesiGNN denotes the full pipeline, including model refinement in Section 3.3.
-Our experiments use NAS-Bench-Graph [35] for accumulated knowledge and a unified search space.
-To prevent data leakage, when a benchmark dataset is treated as unseen, it is completely excluded from the knowledge retrieval process and anonymized, ensuring fair evaluation.
-Out-of-distribution datasets (DBLP, Flickr, Actor) have access to all benchmark graphs as knowledge sources.
-We have validated that LLMs, given only dataset descriptions, lack prior knowledge of effective architectures in the NAS-Bench-Graph search space in Appendix D.1.
-Hyperparameters and more details are in Appendix B.1.
-4.2 Main Results
-Table 2: Initial performance comparison (accuracy) of GNNs designed by different methods. $N_{s}=3$ for Desi-Init. Best overall results are in bold, and best among each category are underlined.
+Using the computed task similarity set $\mathcal{S}^{u}$ for $G^{u}$, we rank the benchmark graphs and select the Top-$N_{s}$ most similar sources.
+The knowledge pool $\mathcal{K}$ for $G^{u}$ is then constructed by aggregating the top-performing models from them:
+(6) $\mathcal{K}=\bigcup_{G^{i}\in\text{Top-}{N_{s}}(\mathcal{S}^{u})}\{(G^{i},\{\theta_{im}^{*}\}_{m=1}^{N_{m}})\},$
+The top-performing models $\{\theta_{i}^{*}\}$ act as a truncated prior $\mathbb{P}(\theta_{i}^{*})$, summarizing accumulated design knowledge from benchmark $G^{i}$, while task similarity $\mathcal{S}(G^{u},G^{i})$ serves as a gating module for solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$, indicating the effectiveness of $\theta_{i}^{*}$ on unseen graph $G^{u}$.
+Table 1. Initial performance comparison (accuracy) of GNNs designed by different methods. $N_{s}=3$ for Desi-Init. Best overall results are in bold, and best among each category are underlined.
 Type Model Cora Cit. Pub. CS Phy. Pho. Com. arX. DBL. Fli. Act.
@@
 DesiGNN 81.77(0.40) 71.00(0.09) 77.57(0.29) 90.51(0.42) 92.61(0.00) 92.38(0.06) 84.08(0.66) 72.02(0.18) 85.89(0.21) 55.44(0.06) 37.57(0.62)
+3.3. Model Suggestion and Refinement
+Overall, our formulation of Knowledge Pool $\mathcal{K}$ approximates the posterior $\mathbb{P}(\theta_{i}^{*}|\mathcal{F}(G^{u}))$, prioritizing models that align with data-aware evidence.
+This Bayesian-inspired process ensures solid use of constructed meta-knowledge for designing models in both one-shot and a few attempts.
+Initial Model Suggestion.
+By supplying contextual meta-knowledge of high similarities to the unseen graph $G^{u}$ from $\mathcal{K}$, we leverage LLM as a surrogate meta-controller to sample $N_{s}$ initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ based on the approximated posterior, where $\mathcal{K}_{i}$ represents the specific knowledge from the $i$-th similar benchmark graph in $\mathcal{K}$: $\theta_{ui}\leftarrow\mathcal{LLM}_{IMS}(\mathcal{F}(G^{u},\bar{I}),\mathcal{K}_{i}).$.
+Compared to existing AutoGNNs, our method avoids any model training on unseen data, delivering promising designs by combining fine-grained literature and empirical insights for immediate model deployment.
+Knowledge-Driven Refinement.
+To further optimize initial model proposals, we design a structured, knowledge-driven refinement strategy inspired by expert-like exploration-exploitation trade-offs, as illustrated in Figure 2.
+Our local search strategy iteratively improves models using the constructed meta-knowledge of high similarities to the unseen from $\mathcal{K}$.
+(0) Re-Ranking: Initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ are re-ranked based on their validation performance, and the best-performing proposal $\theta_{u1}$ serves as the refinement starting point.
+(1) Controlled Exploration: Configurations from other models in $\mathcal{K}$ are crossovered with $\theta_{u1}$ to generate $N_{c}$ new candidates.
+(2) Model Promotion Mechanism: Candidates are ranked based on their retrieved performances on the benchmark dataset of $\mathcal{K}_{1}$, with the most promising candidate $\theta_{u}^{t^{\prime}}$ advanced for further refinement at iteration $t$.
+(3) Directional Exploitation: The LLM meta-controller $\mathcal{LLM}_{KDR}$ mutates $\theta_{u}^{t^{\prime}}$ using user requirements, task descriptions, search space details, and previous training logs.
+(4) Evaluation and Update: Each refined candidate $\theta_{u}^{t}$ is validated and added to the optimization trajectory $\Theta^{T}$.
+The best-performing model is updated as $\theta_{u}^{*}$.
+This refinement process balances exploration (testing diverse configurations) and exploitation (focusing on high-potential candidates), leveraging prior knowledge in the textual space to efficiently optimize model designs.
+4. Experiments
+4.1. Experimental Settings
+Task and Datasets.
+We conduct node classification studies using 8 benchmark datasets from NAS-Bench-Graph (Qin et al., 2022): Cora, Citeseer, PubMed (Sen et al., 2008), CS, Physics, Photo, Computer (Shchur et al., 2018), and ogbn-arXiv (Hu et al., 2020).
+We include 3 extra datasets for out-of-distribution (e.g., heterophily) generalization: DBLP (Bojchevski and Günnemann, 2017), Flickr (Zeng et al., 2019), and Actor (Pei et al., 2020).
+Baselines.
+We compare DesiGNN with 4 categories of 17 baselines: (1) Manually Designed GNNs: 7 GNNs (Kipf and Welling, 2016; Veličković et al., 2017; Hamilton et al., 2017; Xu et al., 2018; Defferrard et al., 2016; Bianchi et al., 2021; Morris et al., 2019) showcasing human experts’ efforts in designing tailored models for specific types of graphs; (2) Classical Automated Approaches: 3 typical automated search strategies (Li and Talwalkar, 2020; Real et al., 2019; Zoph and Le, 2016), plus 2 advanced AutoGNN systems (Gao et al., 2021; Zhou et al., 2022), all operating within the same search space as DesiGNN for fairness; (3) Meta-based Approaches: 3 meta-based, semi-supervised methods (You et al., 2020; Qin et al., 2022; Cao et al., 2023) that recommend models from the most similar benchmark graphs or based on meta-knowledge after testing anchor models on the unseen task; (4) LLM-Based Approaches: 2 recent SOTA LLM-Based methods (Wang et al., 2023a; Dong et al., 2023) that leverage LLMs for iterative architecture search.
+Evaluation Metrics.
+To ensure reliability, we report the average accuracy and standard deviation over 10 runs.
+For automated methods, we also analyze the best-so-far accuracy after validating 1-30 model proposals to measure short-run efficiency.
+To assess efficiency under varying LLM API traffic, we set the basic time unit as the time to validate a model proposal on the unseen dataset.
+Implementation Details.
+Our code [^1] Full code, data, prompts: https://github.com/jilwang84/DesiGNN.
+is implemented in PyTorch and LangChain (GPT-4).
+For clarity, Desi-Init refers to initial model proposals, while DesiGNN denotes the whole pipeline, including model refinement.
+We use NAS-Bench-Graph (Qin et al., 2022) to accumulate knowledge and provide a unified search space.
+To prevent data leakage, when a benchmark dataset is treated as unseen, it is completely excluded from the knowledge retrieval process and anonymized to ensure fair evaluation.
+We have validated that LLMs, given only dataset descriptions, lack prior knowledge of effective architectures in the NAS-Bench-Graph search space.
+Figure 5. Short-run performance of DesiGNN compared to stronger automated baselines.
+4.2. Main Results
 Initial Model Suggestions.
-First, we use Desi-Init in Table 2 to refer to the initial model suggestion part in Section 3.3, i.e., the first GNN suggested by LLM within seconds.
+First, we use Desi-Init in Table 1 to refer to the initial model suggestion part in Section 3.3, i.e., the first GNN suggested by LLM within seconds (avg.
+$0.07$ sec of DesiGNN system time cost + avg.
+$10.77$ sec each for LLM API communication, which may vary among users and API providers).
 While human-designed GNNs like GCN and GAT excel on popular datasets (e.g., Cora, Citeseer), they fail to generalize to less common heterophily datasets such as Flickr and Actor.
-In contrast, Desi-Init surpasses these overly specialized models, demonstrating robustness across diverse graph datasets.
-Notably, Desi-Init outperforms semi-supervised meta-based baselines (Kendall, Overlap, AutoTransfer), validating our contributions of establishing a proficient “graph-GNN-performance” meta-knowledge through unsupervised task similarity assessment and graph understanding with adaptive properties.
-Furthermore, compared to automated baselines, Desi-Init achieves superior performance across all 11 datasets, underscoring the strength of our knowledge-driven approach over conventional automated methods (lack domain-specific insights) and LLM-based methods (rely only on commonplace knowledge).
-Besides, DesiGNN’s initial model proposals rank in the top 5.77% of all possible architectures, as shown in Table 7 at Appendix B.2.
-These results confirm that Desi-Init provides high-quality, tailored model proposals with negligible computational cost by capturing and applying domain-specific knowledge effectively, i.e., alleviating the expertise and efficiency challenges.
-Table 3: Computational resources (in #proposal validations) needed by baselines to reach DesiGNN’s 10-validation performance.
-means failing to achieve within 100 validations in at least 1 of 10 trials.
+In contrast, Desi-Init surpasses these overly specialized models, demonstrating robustness across diverse graphs.
+Notably, Desi-Init outperforms semi-supervised meta-based baselines, validating our contributions of establishing proficient “graph-GNN-performance” meta-knowledge through unsupervised task similarity assessment and graph understanding with adaptive properties.
+Furthermore, compared to automated baselines, Desi-Init achieves superior performance across all 11 datasets, underscoring the strength of our knowledge-driven approach over conventional automated methods (which lack domain-specific insights) and LLM-based methods (which rely solely on commonplace knowledge).
+Overall, DesiGNN’s initial model proposals rank in the top 5.77% of all possible architectures.
+Table 2. Computational resources (in #proposal validations) needed by baselines to reach DesiGNN’s 10-validation performance. contains failed cases within 100 validations.
 #Model Proposal Validations
@@
 GHG. 20-30 $\infty$ 40-50 10 $\infty$ 20-30 20-30 30-40 $\infty$ $\infty$
-Model Refinement and Short-run Efficiency.
-Building on the initial proposals, DesiGNN (full pipeline) can rapidly refine models to outperform all manually designed and automated GNNs across 11 datasets, demonstrating its efficient use of meta-knowledge to exploit the high potential range.
-Figure 4 highlights the short-run efficiency of DesiGNN, which achieves significant performance improvements with fewer model proposal validations compared to automated baselines (full results are presented in Table 8, 9, 10 and Figure 8, 9 at Appendix B.2).
-Unlike those methods requiring extensive validations to identify feasible models, DesiGNN emulates human-like refinement strategies, reducing computational costs and accelerating optimization.
-Specifically, Table 3 shows that other automated methods require significantly more model validations to match DesiGNN’s performance after just 10 validations.
-These results confirm that DesiGNN effectively addresses computational inefficiencies by leveraging accumulated knowledge to enable immediate feedback and rapid optimization, delivering consistently superior performance with minimal computational costs or under limited budgets.
-Figure 4: Short-run performance of DesiGNN compared to stronger automated baselines.
-Figure 5: Task similarities computed by empirical performance (target), two meta-based baselines, and ours.
-Kendall’s $\tau$ correlation coefficient with the empirical ground truth is shown in parentheses.
-4.3 Effective Knowledge via Task Similarity
-To assess whether our unsupervised task similarity in Equation 5 reflects the probability of successful knowledge transfer defined in Definition 1, we quantify the hit rate, defined as the tighter probability $\delta^{*}$ that the empirically most relevant benchmarks (i.e., those maximizing ($\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$) are included in the Top-$N_{s}$ benchmarks retrieved by $\mathcal{S}$ ($N_{s}$ is the tolerance).
-A higher hit rate directly
-Table 4: Hit rate $\delta^{*}$ across unseen datasets.
-Top-s Benchs $N_{s}$=1 $N_{s}$=2 $N_{s}$=3
-Kendall 0 0.25 0.625
-Overlap 0 0.375 0.625
-Property Similar. (w/o LLMs) 0.375 0.5 0.625
-Description Only (with LLMs) 0.262 0.5 0.575
-Property Only (with LLMs) 0.375 0.6 0.725
-Both (with LLMs) 0.275 0.5 0.6
-supports the probabilistic claim in Definition 1, where $\mathcal{S}$ correlates with the likelihood of transferring high-performing models from $G^{i}$ to $G^{u}$.
-As shown in Table 4, our unsupervised task similarity based on graph properties achieves the highest hit rate across all tolerance settings, outperforming methods relying on Kendall or Top 5% similarities and also the purely statistical property similarity (without LLMs to conduct adaptive knowledge elicitation and alignment).
-This demonstrates that our performance-oriented $\mathcal{S}$, derived from key graph properties and adaptive weighting by LLMs, effectively captures the meta-knowledge required for data-aware model design.
-Besides, we observed that there is a negative effect when incorporating descriptive inputs in similarity assessment, suggesting that some semantic properties (e.g., being a citation graph) are less relevant in designing model architecture (case study in Figure 1(b) and Appendix D.2).
-We further visualize the pairwise similarities between graphs in Figure 3(b)5, showing that our unsupervised task similarity closely aligns with empirical performance similarity (the highest Kendall’s $\tau$ correlation).
-This further validates the claim that $\mathcal{S}$ approximates a pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$ well, reinforcing its role in supporting the Bayesian-inspired framework in Section 3.2.
-More Results.
-Due to limited space, the ablation study on main experiments, graph understanding, model suggestion, and model refinement are presented in Appendix C.1, C.2, C.3, C.4, respectively.
-Various LLMs’ innate knowledge, reasoning capabilities, and prompts are studied in Appendix D.
-5 Conclusion
-We present DesiGNN, a knowledge-centered framework that demonstrates deep expertise for proficiently designing data-aware GNNs and eliminates the external noise by systematically converting accumulated model design experiences into fine-grained meta-knowledge well fitted to meta-learning with LLMs.
-Leveraging a performance-oriented task similarity grounded in empirically validated graph properties and adaptive knowledge elicitation via LLMs, DesiGNN bridges the gap between literature insights and practical, data-specific model deployment.
-DesiGNN integrates graph understanding, effective knowledge retrieval, and strategic model suggestion and refinement following a Bayesian perspective, enabling tailored GNN designs for diverse graphs in one shot or a few attempts.
-Extensive experiments demonstrate that DesiGNN expedites the design process by not only delivering top-tier GNN designs within seconds but also achieving superior performance with minimal costs.
-Societal Impact
-Our work contributes to the growing movement toward high-level automation in artificial intelligence by enabling the proficient, data-aware design of Graph Neural Networks (GNNs) through structured knowledge reuse and language model alignment.
-As AI systems become more complex and specialized, the ability to automate model design across unseen tasks without human expertise offers transformative potential for a wide range of real-world applications.
-DesiGNN lowers the barrier to entry for applying advanced graph-based machine learning, making it more accessible to scientists, engineers, and practitioners in domains such as biology, social networks, and recommendation systems.
-By replacing ad hoc manual tuning with systematic and interpretable design strategies, our framework may promote reproducibility, consistency, and scalability in the deployment of machine learning systems.
-More broadly, this work aligns with the vision of LLM-powered agents acting as autonomous scientific assistants—able to rapidly translate raw data into effective models—thereby accelerating innovation in scientific discovery, infrastructure optimization, and decision-making.
-We hope our contribution supports the responsible expansion of AI automation by enabling efficient and expert-level design processes at scale.
-However, such high-level automation also warrants caution.
-While DesiGNN significantly reduces computational cost and expertise requirements, its efficacy may be affected by the quality and representativeness of the knowledge source, along with the reliability of LLMs’ inherent contextual reasoning capability.
-Additionally, while our method operates a level removed from direct applications, any downstream deployment of GNNs—e.g., in social network analysis or recommender systems—can have far-reaching societal implications such as privacy leakage or algorithmic bias if not handled responsibly.
-Limitations and Future Work
-While the current implementation focuses on node classification and leverages a fixed search space from open benchmarks, future work can develop more nuanced mappings between more diverse datasets, more delicate configuration spaces, and the corresponding performance on other tasks, e.g., link prediction.
-Notably, it is worthwhile to study how to select or synthesize benchmark datasets that are representative of their top-performing model patterns so that the knowledge base can maintain maximum knowledge efficacy and better mitigate the out-of-distribution challenges.
-Besides, our research focuses on enhancing the short-run effectiveness of automated GNN designs, leaving long-run knowledge-driven optimization with a more nuanced search tradeoff as an appealing direction.
-Another potential is to extend the currently finite search space to an unlimited design space.
-Looking ahead, there are also promising avenues for extending our knowledge-driven logic to broader machine-learning contexts beyond graph data, e.g., image and tabular data.
-Appendix A Methodology
-In this section, we provide a detailed explanation of the methodology behind our DesiGNN framework, including Graph Understanding (Section 3.1), Knowledge Retrieval (Section 3.2), and GNN Model Suggestion and Refinement Modules (Section 3.3).
-A.1 Algorithm
-The DesiGNN methodology is detailed in Algorithm 1.
-The algorithm is divided into three main phases: Graph Understanding, Knowledge Retrieval, and GNN Model Suggestion and Refinement.
-It also includes details on the Self-evaluation Mechanism, Empirical Filtering Mechanism, Adaptive Filtering Mechanism, Re-rank Mechanism, Controlled Exploration, Model Promotion Mechanism, Directional Exploitation, and Optimization.
-(a) Property confidences across graphs.
-(b) Average confidences of graph properties.
-Figure 6: Empirical studies of graph property confidence as a stronger indicator of effective meta-knowledge in GNNs design.
-Algorithm 1 DesiGNN Methodology
-Input: Unseen Graph Dataset $G^{u}$
-Data: Benchmark Datasets $\{G^{i}\}_{i=1}^{n}$, Benchmark Knowledge Base $\mathcal{H}_{G}:\Theta\to P$, Self-evaluation Bank $BE$, Task-aware LLMs $\mathcal{LLM}_{GDC}$, $\mathcal{LLM}_{IMS}$, $\mathcal{LLM}_{KDR}$, Empirical Filter $\mathcal{F}(\cdot,\cdot)$, #Properties $N_{f}$, Knowledge Pool size $N_{s}$, #Top models $N_{m}$, #Candidates $N_{c}$, stop criteria $maxTrials$
-Output: Optimized GNN architecture $\theta^{*}_{u}$ and model parameters $\omega^{*}$
-1:  Phase 1: Graph Understanding Module
-2:  if $\mathcal{F}(\cdot,\cdot)$ needs initialization or update then
-3:     Self-evaluation Mechanism
-4:     for each anchor dataset $G^{i}$ and property $g_{k}\in BE$ do
-5:        for each other dataset $G^{j}\in BE,G^{j}\neq G^{i}$ do
-6:           Compute $d_{k}^{ij}=|g_{k}^{i}-g_{k}^{j}|$ and retrieve $p^{ij}\in BE$
-7:        end for
-8:        $\mathcal{SR}(i,k)=\text{argsort}(\{d_{k}^{ij}\mid j\neq i,G^{j}\in\mathbf{G}\},\text{ASC})$
-9:        $\mathcal{ER}(i)=\text{argsort}(\{p^{ij}\mid j\neq i,G^{j}\in\mathbf{G}\},\text{DESC})$
-10:        $I(G^{i},g_{k}):=\text{KendallCorr}\big{(}\mathcal{SR}(i,k),\mathcal{ER}(i)\big{)}$
-11:     end for
-12:     $\bar{I}(g_{k})=\frac{1}{n}\sum_{i=1}^{n}I(G^{i},g_{k})$
-13:     Initialize/Update $\mathcal{F}(\cdot,\cdot)=\text{Top}_{N_{f}}(\{\bar{I}(g_{k})\})$
-14:  end if
-15:  Generate unseen dataset’s description with $\mathcal{F}(G^{u},\bar{I})$ and benchmark datasets’ descriptions with $\mathcal{F}(G^{i},\bar{I})$
-16:  Phase 2: Knowledge Retrieval Module
-17:  Compute $\mathcal{S}^{u}=\{\frac{1}{N_{f}}\sum_{k=1}^{N_{f}}w^{u}_{k}\cdot\frac{\bar{I}(g_{k})}{1+\hat{d_{k}^{ui}}}\}_{i=1}^{n}$; $w^{u}_{k}$ is determined by $\mathcal{LLM}_{GDC}\big{(}\mathcal{F}(G^{u},\bar{I}),\mathcal{F}(\{G^{i}\}_{i=1}^{n},\bar{I})\big{)}$
-18:  Retrieve $\mathcal{K}=\bigcup_{G^{i}\in\text{Top-}{N_{s}}(\mathcal{S}^{u})}\{(G^{i},\{\theta_{im}^{*}\}_{m=1}^{N_{m}})\}$
-19:  Phase 3: Initial Model Suggestion
-20:  for each $\mathcal{K}_{i}\in\mathcal{K}$ do
-21:     $\theta_{ui}\leftarrow\mathcal{LLM}_{IMS}(\mathcal{F}(G^{u},\bar{I}),\mathcal{K}_{i})$
-22:     $p^{ui}=\mathcal{H}(\theta_{ui},\omega;G^{u})$
-23:  end for
-24:  $BE\leftarrow BE\cup\{(G^{u},\{g_{k}^{u}\},\{p^{ui}\})\}$
-25:  Phase 4: Model Proposal Refinement
-26:  $R=\text{argsort}(\{p_{ui}\},\text{DESC})$
-27:  $\mathcal{K}_{1:N_{s}}=\{(G^{R[j]},\{\theta_{R[j]m}^{*}\}_{m=1}^{N_{m}})\}_{j=1}^{N_{s}}$
-28:  $\Theta_{u}^{T}\leftarrow(\theta_{u}^{R[1]},p_{u}^{R[1]})$; $\theta^{*}_{u}=\theta_{u}^{T}[1]$
-29:  for $t=N_{s}$ to $maxTrials$ do
-30:     $C^{t}=\{\theta_{ui}^{t^{\prime}}\mid\text{Crossover}(\theta^{*}_{u},\mathcal{K}_{2:N_{s}})\}_{i=1}^{N_{c}}$
-31:     $\theta_{u}^{t^{\prime}}=\text{argsort}(\{\mathcal{H}_{\mathcal{K}_{1}}(\theta_{ui}^{t^{\prime}})\mid\theta_{ui}^{t^{\prime}}\in C^{t}\},\text{DESC})[1]$
-32:     $\theta_{u}^{t}\leftarrow\mathcal{LLM}_{MPR}(\theta_{u}^{t^{\prime}},\Theta^{T},\mathcal{K}_{1})$
-33:     $p_{u}^{t}=\mathcal{H}(\theta_{u}^{t},\omega;G^{u})$
-34:     $\Theta_{u}^{T}\leftarrow(\theta_{u}^{t},p_{u}^{t})$
-35:     if $p_{u}^{t}>p_{u}^{*}$ then
-36:        $\theta^{*}_{u},\omega^{*}=\theta_{u}^{t},\omega$
-37:        $p_{u}^{*}=p_{u}^{t}$
-38:     end if
-39:  end for
-40:  return $\theta^{*}_{u},\omega^{*}$
-A.2 Graph Understanding Module
-The Graph Understanding Module (Section 3.1) is designed to automatically analyze the graph properties to enhance task comprehension beyond mere user input and establish an empirical foundation for our meta-knowledge construction and unsupervised task similarity.
-The module initially prioritizes the most influential properties based on their confidences averaged across all benchmark datasets, which is illustrated in Figure 6(b).
-The selected features are then used to generate textual descriptions of graph datasets, which are subsequently employed to perform adaptive elicitation, align knowledge, and finally identify similar benchmark datasets to retrieve data-aware model design knowledge.
-A.2.1 Graph Properties Pool
-We employ a comprehensive set of $16$ graph properties by analyzing the graph learning literature [54, 47].
-Illustrated in Table 1, the corresponding properties name from $g_{1}$ to $g_{16}$ are listed below (and are also in the initialized order based on their confidences $\bar{I}(g_{k})$): the average clustering coefficient, average betweenness centrality, density, average degree centrality, average closeness centrality, average degree, edge count, graph diameter, average shortest path length, assortativity, average eigenvector centrality, feature dimensionality, node count, node feature diversity, connected components, and label homophily.
-Together, they could capture the topology of graphs, which commonly motivates the specific design of GNNs in research.
-The underlined properties are measured on sampled subgraphs to reduce computational overhead and ensure scalability.
-A.2.2 Statistical and Empirical Similarities
-To empirically study the missing meta-level understanding about graph-GNN-performance, we start with 16 graph properties $\{g_{k}\}_{k=1}^{16}$ from the literature to infer graph similarity from two aspects: 1) the statistical distance ranking based on graph properties $g_{k}$, and 2) the empirical performance ranking of transferred top-performing GNNs from other graphs $\{G^{j}\mid j\neq i\}$ to anchor graph $G^{i}$.
-Specifically, given two graphs $G^{i}$ and $G^{j}\in\mathbf{G}$, we first compute the statistical distance $d_{k}^{ij}=|g_{k}^{i}-g_{k}^{j}|$ between $G^{i}$ and $G^{j}$ based on $g_{k}$, i.e., the reverse ranking of $\{d_{k}^{ij}\}_{k=1}^{16}$ can assess the sectional topological similarities between graphs $G^{i}$ and $G^{j}$ after normalization.
-Then, to quantify empirical transferability, we let $p^{ij}$ denote the performance of transferring the top-performing GNN patterns summarized by LLMs from another graph $G^{j}$ to $G^{i}$ (details in Section 3.3).
-Based on $d_{k}^{ij}$ and $p^{ij}$, we define two similarity rankings of other graphs $\{G^{j}\mid j\neq i,G^{j}\in\mathbf{G}\}$ to graph $G^{i}$:
-$\mathcal{SR}(i,k)$ $=\operatorname{argsort}\left(\{d_{k}^{ij}\mid j\neq i,\,G^{j}\in\mathbf{G}\},\,\text{ASC}\right),$ (7)
-$\mathcal{ER}(i)$ $=\operatorname{argsort}\left(\{p^{ij}\mid j\neq i,\,G^{j}\in\mathbf{G}\},\,\text{DESC}\right)$ (8)
-where $\mathcal{SR}(i,k)$ represents the statistical similarity ranking based on property $g_{k}$, and $\mathcal{ER}(i)$ represents the empirical ground truth of knowledge transferability when meta-learning with LLMs.
-A.3 Knowledge Retrieval Module
-As the knowledge source of our Knowledge Retrieval Module (Section 3.2), NAS-Bench-Graph [35] is a comprehensive benchmark designed to facilitate unified, reproducible, and efficient evaluations of graph neural architecture search (GraphNAS) methods.
-This benchmark encapsulates a well-defined search space and a rigorous evaluation protocol, which encompasses the training and testing of 26,206 unique graph neural network (GNN) architectures across nine diverse node classification datasets.
-A.3.1 Benchmark Datasets
-The benchmark utilizes nine graph datasets of varied sizes and types, including citation networks (Cora, Citeseer, PubMed) [37], co-authorship graphs (Coauthor CS and Coauthor Physics) [38], co-purchase networks (Amazon Computers and Amazon Photo) [38], and large-scale graphs like ogbn-arXiv [20] from the Open Graph Benchmark.
-These datasets are employed with fixed semi-supervised [35] splits to ensure consistent evaluation settings.
-A.3.2 Search Space Design
-The search space in NAS-Bench-Graph is constructed as a directed acyclic graph (DAG), representing various GNN architectures.
-It includes macro architectural choices constrained to 9 distinct patterns: [0,0,0,0], [0,0,0,1], [0,0,1,1], [0,0,1,2], [0,0,1,3], [0,1,1,1], [0,1,1,2], [0,1,2,2], and [0,1,2,3].
-The operations include seven prominent GNN layer types—GCN [24], GAT [42], GraphSAGE [18], GIN [52], ChebNet [10], ARMA [2], and k-GNN [32]—plus Identity and fully connected layers for residual connections and non-graph structure use, respectively.
-Figure 7: Conceptual illustration of controlled exploration and directional exploitation in GNN model suggestion and refinement.
-Table 5: Statistics of the datasets.
-Dataset #Vtx #Edges Feat. Cls Metric
-Cora 2,708 5,429 1,433 7 Acc.
-Citeseer 3,327 4,732 3,703 6 Acc.
-PubMed 19,717 44,338 500 3 Acc.
-CS 18,333 81,894 6,805 15 Acc.
-Physics 34,493 247,962 8,415 5 Acc.
-Photo 7,487 119,043 745 8 Acc.
-Computers 13,381 245,778 767 10 Acc.
-arxiv 169,343 1,166,243 128 40 Acc.
-proteins 132,534 39,561,252 8 112 ROC
-DBLP 17,716 105,734 1,639 4 Acc.
-Flickr 89,250 899,756 500 7 Acc.
-Actor 7,600 30,019 932 5 Acc.
-A.4 GNN Model Suggestion and Refinement
-As introduced in Section 3.3 and illustrated in Figure 2, our model design refinement cycle continues iteratively, with each step methodically enhancing the model architecture based on accumulated knowledge and feedback.
-This process integrates user requirements, retrieved top-performing models from $\mathcal{K}$, and search space description (e.g., macro architectures and operations).
-The detailed procedure is as follows: (0) Re-Ranking: Initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ are re-ranked based on their validation performance, and the best-performing proposal $\theta_{u1}$ serves as the refinement starting point.
-(1) Controlled Exploration: Configurations from other models in $\mathcal{K}$ are crossovered with $\theta_{u1}$ to generate $N_{c}$ new candidates.
-(2) Model Promotion Mechanism: Candidates are ranked based on their retrieved performances on the benchmark dataset of $\mathcal{K}_{1}$, with the most promising candidate $\theta_{u}^{t^{\prime}}$ advanced for further refinement at iteration $t$.
-(3) Directional Exploitation: The LLM meta-controller $\mathcal{LLM}_{KDR}$ mutates $\theta_{u}^{t^{\prime}}$ using user requirements, task descriptions, search space details, and previous training logs.
-(4) Evaluation and Update: Each refined candidate $\theta_{u}^{t}$ is validated and added to the optimization trajectory $\Theta^{T}$.
-The best-performing model is updated as $\theta_{u}^{*}$.
-As conceptually illustrated in Figure 7, the search space is navigated in a controlled manner through simulated crossovers between $\theta^{*}$ and configurations from $\mathcal{K}_{2:N_{s}}$, ensuring that exploration remains within high-potential range.
-Among these generated candidates, the one that demonstrates the highest potential based on its performance on the closest benchmark dataset is identified as the most promising for the unseen dataset.
-This candidate is then subjected to further exploitation by the LLM, guided by the most valuable directional insights derived from $\mathcal{K}_{1}$.
-Our empirical results support the short-run effectiveness of this local search strategy in quickly refining models.
-Our retrieval-then-verified promotion strategy not only ensures a reliable and efficient selection of new proposals by leveraging pre-existing, empirically validated knowledge but also requires only a single evaluation per optimization step.
-This sharply contrasts with existing traditional and LLM-based methods that necessitate accumulating knowledge from scratch and running multiple evaluations in parallel, thereby reducing computational overhead and accelerating the refinement process.
-Appendix B Experiments
-B.1 Experimental Settings
-Table 6: Summary of hyperparameters used for each dataset
-Dataset #Pre-process #Post-process Dimension Dropout Optimizer Learning Rate Weight Decay Epochs
-Cora 0 1 256 0.7 SGD 0.1 0.0005 400
-Citeseer 0 1 256 0.7 SGD 0.2 0.0005 400
-PubMed 0 0 128 0.3 SGD 0.2 0.0005 500
-CS 1 0 128 0.6 SGD 0.5 0.0005 400
-Physics 1 1 256 0.4 SGD 0.01 0 200
-Photo 1 0 128 0.7 Adam 0.0002 0.0005 500
-Computers 1 1 64 0.1 Adam 0.005 0.0005 500
-ogbn-arxiv 0 1 128 0.2 Adam 0.002 0 500
-ogbn-proteins 1 1 256 0 Adam 0.01 0.0005 500
-DBLP 1 1 256 0.5 SGD 0.1 0.0005 300
-Flickr 1 1 128 0.5 Adam 0.001 0.0005 300
-Actor 1 1 128 0.5 Adam 0.005 0.0005 400
-Table 7: Performance Ranking of Initial Model Proposals.
-Performance Rank (%)
-Avg. Cora Cit. Pub. CS Phy. Pho. Com arX.
-5.765 5.899 3.324 4.926 11.036 0.053 3.144 7.353 10.383
-B.1.1 Task and Datasets.
-We evaluate our DesiGNN framework on eleven diverse graph datasets, including eight out of nine benchmark datasets { Cora [37], Citeseer [37], PubMed [37], CS [38], Physics [38], Photo [38], Computer [38], ogbn-arXiv [20] } from NAS-Bench-Graph [35] and three additional datasets { DBLP [3], Flickr [56], Actor [34] }.
-All datasets are used to conduct node classification tasks, and their statistics are summarized in Table 7.
-The reason for excluding ogbn-proteins [20] in evaluation is that NAS-Bench-Graph only recorded partial model performance mapping due to explosion and out-of-memory errors, making it hard to conduct benchmarking comparisons.
-B.1.2 Training Hyperparameters
-The training hyperparameters used in different datasets are summarized in Table 6.
-For the nine benchmark datasets, we use the same hyperparameters as the NAS-Bench-Graph [35] benchmark to ensure fair comparisons.
-These hyperparameters were selected based on random searches on 30 anchor GNN architectures.
-The hyperparameters for the three additional datasets are recommended by LLMs following the same procedure from Graph Understanding (Section 3.1) to Initial Model Suggestion (Section 3.3), except the retrieved model architecture knowledge (Section 3.2) is replaced by the tuned hyperparameters of the corresponding benchmark datasets [35].
-All the hyperparameters are fixed across all experiments for fairness.
-All experiments are conducted on a single NVIDIA RTX 3080 GPU and any GPU that could run a conventional GNN training can work.
-Notably, DesiGNN’s initial model recommendation can be delivered without GPU.
-Table 8: Best model performance (accuracy) after 10 model validations on the benchmark datasets.
-ACC (STD) %
-Type Model Cora Citeseer PubMed CS Physics Photo Computer arXiv DBLP Flickr Actor
-Auto. AutoGL GNAS 80.89(0.35) 67.45(1.40) 76.36(0.50) 88.94(0.93) 89.55(2.17) 87.76(3.37) 75.70(5.54) 69.84(1.46) 84.67(0.29) 54.69(1.02) 35.12(1.71)
-Auto-GNN 80.63(1.13) 68.45(0.72) 76.35(0.66) 88.46(2.52) 90.94(0.48) 90.08(1.63) 78.39(3.76) 67.78(1.91) 84.84(0.36) 51.27(2.01) 31.32(5.31)
-Auto. NNI Random 80.79(0.60) 65.53(4.95) 75.79(0.79) 88.95(1.41) 90.68(0.27) 90.76(1.00) 76.95(3.00) 71.27(0.52) 82.68(5.89) 52.93(1.21) 33.33(3.35)
-EA 81.27(0.59) 67.29(1.37) 74.92(0.88) 88.08(1.61) 91.05(0.64) 89.63(1.26) 80.46(1.88) 70.40(0.50) 85.15(0.33) 54.38(0.87) 33.83(2.65)
-RL 80.28(0.77) 67.39(1.06) 75.39(1.34) 88.47(1.53) 90.70(0.90) 88.84(0.39) 77.48(3.90) 69.12(1.78) 84.46(0.23) 53.34(0.94) 33.61(2.43)
-LLM -based GPT4GNAS 80.33(0.39) 69.37(0.01) 76.40(0.20) 90.24(0.06) 91.06(0.61) 91.51(0.78) 82.88(1.25) 70.85(0.22) 85.45(0.16) 55.12(0.16) 36.14(0.54)
-GHGNAS 80.51(0.45) 69.30(0.22) 76.49(0.32) 90.28(0.29) 90.94(0.28) 91.64(0.47) 82.40(1.21) 70.88(0.38) 85.42(0.16) 55.00(0.18) 36.56(0.49)
-Our DesiGNN 81.69(0.54) 70.58(0.46) 77.17(0.29) 90.22(0.48) 92.61(0.00) 92.25(0.21) 83.15(0.61) 71.92(0.15) 85.84(0.27) 55.20(0.09) 36.58(2.63)
-Table 9: Best model performance (accuracy) after 20 model validations on the benchmark datasets.
-ACC (STD) %
-Type Model Cora Citeseer PubMed CS Physics Photo Computer arXiv DBLP Flickr Actor
-Auto. AutoGL GNAS 81.35(0.23) 69.49(0.65) 76.94(0.21) 90.12(0.27) 91.64(0.52) 91.47(0.19) 83.15(1.23) 71.49(0.43) 85.51(0.07) 55.22(0.21) 36.82(0.59)
-Auto-GNN 81.07(0.65) 69.78(0.52) 77.21(0.58) 90.36(0.35) 91.92(0.65) 91.52(1.16) 82.73(1.17) 71.50(0.30) 85.65(0.18) 54.99(0.18) 37.10(0.91)
-Auto. NNI Random 80.97(0.45) 69.57(0.23) 76.69(0.15) 90.20(0.27) 91.80(0.31) 91.74(0.46) 83.21(0.61) 71.54(0.23) 85.37(0.13) 55.01(0.34) 37.31(0.58)
-EA 81.27(0.59) 67.82(0.93) 76.33(0.70) 89.53(1.03) 91.38(0.51) 91.67(0.49) 82.16(1.17) 71.48(0.27) 85.64(0.25) 55.01(0.23) 36.73(1.48)
-RL 80.53(0.48) 69.69(0.51) 77.33(0.62) 90.30(0.44) 91.48(0.51) 91.75(0.31) 82.46(0.99) 71.22(0.24) 85.44(0.16) 55.01(0.13) 37.52(0.31)
-LLM -based GPT4GNAS 80.73(0.15) 69.43(0.07) 76.88(0.41) 90.44(0.06) 91.84(0.52) 91.79(0.43) 83.54(0.53) 71.31(0.27) 85.57(0.14) 55.12(0.16) 36.66(0.73)
-GHGNAS 80.66(0.42) 69.44(0.06) 76.93(0.22) 90.42(0.06) 91.59(0.56) 92.03(0.39) 82.93(0.63) 71.44(0.31) 85.43(0.15) 55.06(0.11) 37.01(0.57)
-Our DesiGNN 81.69(0.52) 70.87(0.09) 77.27(0.41) 90.41(0.39) 92.61(0.00) 92.29(0.05) 83.43(0.77) 71.98(0.16) 85.85(0.29) 55.32(0.14) 37.57(0.62)
-Table 10: Best model performance (accuracy) after 30 model validations on the benchmark datasets.
-ACC (STD) %
-Type Model Cora Citeseer PubMed CS Physics Photo Computer arXiv DBLP Flickr Actor
-Auto. AutoGL GNAS 81.35(0.23) 69.67(0.47) 76.97(0.22) 90.18(0.31) 91.92(0.43) 91.61(0.23) 83.23(1.13) 71.54(0.36) 85.58(0.08) 55.23(0.20) 37.23(0.56)
-Auto-GNN 81.46(0.41) 69.78(0.52) 77.37(0.51) 90.50(0.20) 91.92(0.65) 91.97(0.40) 83.06(0.91) 71.61(0.35) 85.66(0.17) 55.20(0.15) 37.50(0.83)
-Auto. NNI Random 81.17(0.44) 69.63(0.16) 77.25(0.44) 90.39(0.14) 91.81(0.31) 92.04(0.32) 83.62(0.63) 71.55(0.22) 85.48(0.15) 55.17(0.12) 37.39(0.59)
-EA 81.27(0.59) 68.05(0.70) 76.51(0.83) 89.70(0.80) 91.60(0.65) 91.78(0.60) 83.00(0.84) 71.62(0.26) 85.66(0.24) 55.09(0.21) 37.44(1.86)
-RL 80.91(0.24) 69.88(0.24) 77.33(0.62) 90.47(0.19) 91.83(0.38) 91.75(0.31) 82.48(0.96) 71.47(0.08) 85.52(0.23) 55.05(0.13) 37.52(0.31)
-LLM -based GPT4GNAS 81.31(0.24) 69.43(0.07) 76.90(0.41) 90.44(0.06) 92.12(0.21) 92.21(0.13) 83.96(0.83) 71.67(0.44) 85.57(0.14) 55.12(0.16) 36.70(0.70)
-GHGNAS 81.39(0.07) 69.64(0.38) 76.93(0.22) 90.42(0.06) 92.06(0.17) 92.38(0.02) 83.28(0.43) 71.84(0.03) 85.47(0.11) 55.06(0.11) 37.01(0.57)
-Our DesiGNN 81.77(0.40) 71.00(0.09) 77.57(0.29) 90.51(0.42) 92.61(0.00) 92.38(0.06) 84.08(0.66) 72.02(0.18) 85.89(0.21) 55.44(0.06) 37.57(0.62)
-B.1.3 Module Hyperparameters
-The hyperparameters used in the DesiGNN framework are summarized as follows: the best empirical number of confident graph properties $N_{f}=8$ (without descriptional inputs), the size of the knowledge pool $N_{s}=3$ to balance tolerance with computation cost (consider benchmark datasets with Top-$3$ similarities to the unseen dataset), the number of top models $N_{m}=30$ (only top-performing model design examples), the number of candidates $N_{c}=30$, and the maximum number of trials $maxTrials=30$ for fast deployment scenario.
-These hyperparameters are selected based on the ablation studies in Section C.2,  C.3, and  C.4, which are then fixed across all main experiments.
-B.2 Main Results
-B.2.1 Performance Ranking of Initial Model Proposals
-In this section, Table 7 presents the performance ranking of initial model proposals within the model space defined by NAS-Bench-Graph [35].
-The average performance ranking of initial model proposals across all benchmark datasets is Top-$5.77\%$, indicating that the initial model proposals suggested by our DesiGNN-Init without any prior training are highly competitive.
-B.2.2 Model Refinement and Short-run Efficiency
-In this section, we present the complete results of the short-run experiments in Table 8 (after 10 model validations), Table 9 (after 20 model validations), and Table 10 (after 30 model validations).
-The performance trajectories of all automated baselines are presented in Figure 8 and Figure 9.
-Table 11: The main ablation study on each component in GNN Model Suggestion and Refinement. * is our complete setting.
+Table 3. The main ablation study on each component in GNN Model Suggestion and Refinement. * is our complete setting.
 ACC (STD) %
@@
 w/o Knowledge 80.90 (0.00) 69.93 (0.00) 76.72 (0.00) 90.24 (0.00) 91.82 (0.00) 92.21 (0.00) 83.63 (0.00) 71.71 (0.00)
-Appendix C Ablation Studies
-To delve deeper into the pivotal designs of DesiGNN, we conducted ablation and hyperparameter studies on three key modules, including Graph Understanding (Section 3.1), Knowledge Retrieval (Section 3.2), and GNN Model Suggestion and Refinement (Section 3.3).
-C.1 Main Ablation Study
-We present the main ablation study on the Knowledge Retrieval and the GNN Model Suggestion and Refinement module in Table 11, quantifying the impact of the retrieved knowledge and each mechanism in our framework.
-The results from the initial model suggestion stage underscore the importance of designing a proper Graph Understanding method, demonstrating that relying solely on filtered graph properties in dataset descriptions outperforms any use of descriptional inputs.
-In the short-run performance evaluation after 30 model validations, the Re-rank mechanism significantly improves the knowledge-driven model refinement process when the original order of the Knowledge Pool $\mathcal{K}$ is incorrect.
+Figure 6.
+Task similarities computed by empirical performance (target), two meta-based baselines, and ours.
+Kendall’s $\tau$ correlation coefficient with the empirical ground truth is shown in parentheses.
+Table 4. Hit rate $\delta^{*}$ across unseen datasets.
+Top-s Benchs $N_{s}$=1 $N_{s}$=2 $N_{s}$=3
+Kendall 0 0.25 0.625
+Overlap 0 0.375 0.625
+Property Similar. (w/o LLMs) 0.375 0.5 0.625
+Description Only (with LLMs) 0.262 0.5 0.575
+Property Only (with LLMs) 0.375 0.6 0.725
+Both (with LLMs) 0.275 0.5 0.6
+Model Refinement and Short-run Efficiency.
+Building on the initial proposals, DesiGNN can rapidly refine models to outperform all manually designed and automated GNNs across 11 datasets, demonstrating its efficient use of meta-knowledge to exploit the high potential range.
+Figure 5 highlights DesiGNN’s short-run efficiency, achieving significant performance improvements with fewer model proposal validations than automated baselines.
+Unlike methods that require extensive validation to identify feasible models, DesiGNN emulates human-like refinement strategies, reducing computational costs and accelerating optimization.
+Specifically, Table 2 shows that other automated methods require significantly more model validations to match DesiGNN’s performance after just 10 validations.
+These results confirm that DesiGNN effectively addresses computational inefficiencies by leveraging accumulated knowledge to enable immediate feedback and rapid optimization, delivering consistently superior performance under limited budgets.
+4.3. Effective Knowledge via Task Similarity
+To assess whether our unsupervised task similarity in Equation 5 reflects the probability of successful knowledge transfer defined in Definition 3.1, we quantify the hit rate, defined as the tighter probability $\delta^{*}$ that the empirically most relevant benchmarks (i.e., those maximizing $\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$) are included in the Top-$N_{s}$ benchmarks retrieved by $\mathcal{S}$ ($N_{s}$ is the tolerance).
+A higher hit rate directly supports the probabilistic claim in Definition 3.1, where $\mathcal{S}$ correlates with the likelihood of transferring high-performing models from $G^{i}$ to $G^{u}$.
+As shown in Table 4, our unsupervised task similarity based on graph properties achieves the highest hit rate across all tolerance settings, outperforming methods that rely on Kendall or Top 5% similarities, as well as purely statistical property similarity (without LLMs to conduct adaptive knowledge elicitation and alignment).
+This demonstrates that our performance-oriented $\mathcal{S}$, derived from key graph properties and adaptive weighting by LLMs, effectively captures the meta-knowledge required for data-aware model design.
+Besides, we observed a negative effect when incorporating descriptive inputs into similarity assessment, suggesting that some semantic properties (e.g., being a citation graph) are less relevant to designing the model architecture (see the case study in Figure 1(b)).
+We further visualize the pairwise similarities between graphs in Figures 3(b) and  6, showing that our unsupervised task similarity closely aligns with empirical performance similarity (the highest Kendall’s $\tau$ correlation).
+This further validates the claim that $\mathcal{S}$ approximates a pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$ well, reinforcing its role in supporting the Bayesian-inspired framework in Section 3.2.
+4.4. Main Ablation Study
+We present the main ablation study on the Knowledge Retrieval and the GNN Model Suggestion and Refinement module in Table 3.
+The results from the initial model suggestion stage underscore the importance of designing an effective Graph Understanding method, demonstrating that relying solely on filtered graph properties in dataset descriptions outperforms using descriptional inputs.
+In the short run, after 30 model validations, the Re-rank mechanism significantly improves the knowledge-driven model refinement process when the original order of the Knowledge Pool $\mathcal{K}$ is incorrect.
 Additionally, the Model Promotion mechanism, which simulates the strategy of human experts refining a model based on accumulated knowledge, plays a crucial role in enhancing the efficacy of model refinement.
 Lastly, the Directional Exploration mechanism is empirically beneficial, as it leverages the most relevant model design knowledge and the in-context learning ability of LLMs to further refine the best candidate model promoted.
-Table 12: Average ranks of different combinations across datasets. A higher rank (with 1 being the highest) corresponds to better performance and lower variability in the initial model proposals.
-Combined Rank
-Top-$N_{s}$ Desc. g=0 g=2 g=4 g=6 g=8 g=10 g=16
-$N_{s}$=1 True 2.250 2.312 2.250 2.125 2.437 2.625 2.312
-False 2.562 2.125 2.562 1.812 2.187 2.250 2.125
-$N_{s}$=2 True 2.500 2.187 2.625 1.937 1.687 2.312 2.750
-False 2.625 2.375 2.812 2.000 1.500 2.187 2.250
-$N_{s}$=3 True 2.500 1.937 2.875 1.875 1.937 2.187 2.000
-False 2.125 1.687 2.000 1.750 1.500 1.3750 1.687
-C.2 Ablation Studies on Graph Understanding
-In this section, we study the combined effect of the number of selected properties $N_{f}$ and the usage of descriptional inputs in understanding and comparing graph datasets, under varying sizes $N_{s}$ of the Knowledge Pool, which influences the number of initial proposals suggested by LLMs (Section 3.3).
-The average rankings of different combinations across datasets in Table 12 illustrate that while the optimal $N_{f}$ varies with different $N_{s}$ values, integrating descriptional inputs often diminishes performance, particularly when $N_{f}>6$ and $N_{s}=3$.
-This finding supports our notion of “artificial hallucinations,” where descriptive inputs can hinder the understanding of tasks when sufficient graph properties are provided.
-From another aspect, this result also validates our idea of analyzing graph properties to design tailored GNN models, which came from insights documented in the literature.
-Table 13: The distribution of the best initial model among the Top-3 benchmark.
-Is Best Initial Model (%)
-Scenario Best Second Third
-Descriptional Only 30.0 50.0 20.0
-$N_{f}=6$ 50.0 22.5 27.5
-$N_{f}=8$ 52.5 37.5 10.0
-$N_{f}=10$ 42.5 40.0 17.5
-Table 14: Average performance rank of transferring the Top-$N_{m}$ model designs.
-Average Rank
-Top-$N_{m}$ $N_{m}$=1 $N_{m}$=10 $N_{m}$=20 $N_{m}$=30 $N_{m}$=40 $N_{m}$=50
-2.7077 2.8750 2.8438 2.6875 2.9231 2.8769
-C.3 Ablation Studies on GNN Model Suggestion
-This study extends from Section 3.1 and quantifies the advantage of using the two $N_{s}=2$ or three $N_{s}=3$ most similar benchmarks instead of solely the most similar one $s=1$ (no tolerance).
-Table 13 shows the distribution of the best initial proposal stems from the first, second, or third most similar benchmark.
-For $N_{f}=8$ without descriptive inputs, approximately $53\%$ of trials identify the optimal initial proposal from the most similar benchmark, with only $10\%$ benefiting from the third.
-This data supports a balanced trade-off between performance and efficiency.
-Notably, this study also serves as the motivation for us to design the Re-rank Mechanism in Model Proposal Refinement (Appendix A.4).
-Moreover, we explore the optimal number $N_{m}$ of top-performing examples from each benchmark when establishing the Knowledge Pool and assess the impact of using poor examples.
-Table 14 shows that $N_{m}=30$ yields the best average performance across all dataset combinations ($2160$ records), with further analysis indicating that poor examples do not significantly affect performance outcomes (T-statistic $=-0.1262$, P-value $=0.8998$).
-We thus adopt $N_{m}=30$ without bad examples for an appropriate token length of the LLM prompt.
-C.4 Ablation Studies on GNN Model Refinement
-We investigate the impact of using the $\mathcal{K}_{1}$ Knowledge Pool as context for LLMs and the Re-rank Mechanism that is based on the performance ranking of initial proposals.
-We also study the best number of candidates and the effect of adopting different $N_{s}$ numbers of knowledge bases to guide the controlled search space exploration.
-Table 15 displays T-statistics for various configurations of Knowledge Pool usage and Re-rank Mechanism.
-For the row of w/o candidate, the True/False columns stand for w/o re-rank, and vice versa.
-While statistical significance is not reached (p-value $>0.05$), certain trends emerge:
-  • When $N_{s}=3$ knowledge bases are used for a broader perspective, employing a Re-rank Mechanism consistently adds value.
-This pattern is intuitive as the Re-rank Mechanism will become more influential as the number of knowledge bases increases.
-  • Broader exploration with $N_{s}=3$ and a larger number of candidates (e.g., $30$) is likely effective, as we also have a directional exploitation mechanism to refine the diversity of potentially immature solutions.
-Given the complexity of these patterns, we further assess the performance ranking of GNN Model Refinement hyperparameter combinations across benchmarks to determine the most effective settings.
-Table 16 suggests that the best hyperparameter combination is $N_{s}=3$ and $N_{m}=30$ with both Re-rank Mechanism and knowledge-driven refinement strategy enabled.
-Table 15: T-statistic of the ablation on the usage of knowledge-driven Exploration and Re-rank Mechanism with different candidate number and $N_{s}$.
-Candidate Number= 10 20 30
-w/o Re-rank or w/o Exploration
-Ablation $N_{s}$ True False True False True False
-w/o Exploration 2 -0.280 0.582 0.512 1.977 0.557 0.566
-3 0.863 0.072 0.055 -0.164 -0.054 -0.358
-w/o Re-rank 2 0.647 1.383 -0.334 1.232 -2.092 -0.852
-3 -0.116 -0.808 -0.593 -0.591 -0.751 -1.904
-Table 16: Average performance rank of the GNN Model Refinement hyperparameter combinations across benchmark.
-Candidate Number= 10 20 30
-w/o Re-rank
-$N_{s}$ w/o Exploration True False True False True False
-2 True 7.33 12 7.83 9 13.5 9.17
-False 7.16 14.5 11.7 17.33 16.17 11.83
-3 True 11.7 11.17 11.83 9.5 15.33 7.33
-False 12.33 10.17 11.17 10.17 12.83 7
-Appendix D Case Studies
-D.1 Case Studies: Lack of Prior Knowledge in LLMs
-The first study verifies the inherent knowledge gap illustrated in Figure 1(a): lack of prior knowledge in LLMs about the top-performing models of benchmark datasets within the NAS-Bench-Graph [35] model space (Section A.3).
-The first two columns in Table 20 are cases where the descriptive inputs or graph properties are directly sent to LLMs for model suggestions.
-The results demonstrate that LLMs are only able to suggest commonplace layer connections and operations based on user-provided descriptional inputs (left panel) or key graph properties (middle panel).
-In fact, we have tested all the benchmark datasets in NAS-Bench-Graph [35] with descriptive inputs, and we found only two macro lists and three operation lists that LLM would recommend regardless of the datasets:
-  • Architecture: [0, 1, 2, 3] and [0, 0, 1, 3].
-  • Operations: [’gcn’, ’gat’, ’sage’, ’skip’], [’gcn’, ’gat’, ’sage’, ’gcn’], and [’gcn’, ’gat’, ’gin’, ’sage’],
-After providing the tailored knowledge retrieved from our framework to LLMs (right panel), it can be clearly seen that the model it recommends is specialized and has obvious performance improvements.
-D.2 Case Studies: Artificial Hallucination in LLMs
-The second case study in Table 21 and Figure 1(b) examines the phenomenon of “artificial hallucination”, or the external noise issue, in LLMs when comparing the similarities between the unseen dataset and the benchmark datasets, as detailed in Section 3.2.
-Figure 5 (leftmost) illustrates that the empirically most similar benchmark datasets to PubMed are CS, Physics, and Citeseer, in descending order of similarity.
-When employing our Graph Understanding method in Section 3.1, which leverages confident graph properties (right panel), the top three most similar datasets are Citeseer, CS, and Physics (align well with empirical ground truth).
-However, the current practice that relies solely on descriptive inputs (left panel) identifies Cora, Citeseer, and ogbn-arxiv as the top three, which does not align with the empirical ground truth.
-This discrepancy arises because LLMs overly rely on the shared characteristic of being citation graphs, assuming that citation datasets like PubMed, Cora, Citeseer, and ogbn-arxiv should have similar model design preferences.
-This case study demonstrates that relying solely on external descriptive inputs is insufficient to capture the similarities between datasets, which can overwhelm other pertinent information, leading to inaccurate task similarity assessment and incorrect knowledge retrieval.
-Table 17: Performance comparison of initial model suggestion with different LLMs as meta-controller.
+Table 5. Different LLMs for model suggestion and refinement.
 Dataset Cora Cit. Pub. CS Phy. Pho. Com. arX.
+Initial Model Suggestion
 Llama2 80.26 68.90 75.83 89.84 90.75 91.35 80.27 70.95
 GPT-4 80.31 69.20 76.60 89.64 92.10 91.19 82.20 71.50
-Table 18: Performance comparison of model refinement with different LLMs as meta-controller.
-Dataset Cora Cit. Pub. CS Phy. Pho. Com. arX.
+Model Refinement
 Llama2 80.96 70.38 77.40 90.19 91.47 92.09 83.69 71.63
 GPT-4 81.77 71.00 77.57 90.51 92.61 92.38 84.08 72.02
-Table 19: Performance comparison of different prompt designs on GPT4GNAS.
-Dataset Cora Cit. Pub. CS Phy. Pho. Com. arX.
-Original 81.31 69.43 76.90 90.44 92.12 92.21 83.96 71.67
-Our Prompt 81.38 70.06 77.03 90.22 92.16 92.08 83.78 71.72
-D.3 Case Studies: Different LLMs
-To study the effect of different LLMs as the meta-controller, we adopt an open-source LLM, Llama2-13b, to replace the GPT-4 meta-controller we originally employed in our paper.
-Table 17 shows the comparative results (averaged over five runs) on the initial model proposal, while Table 18 shows the comparative results (averaged over five runs) after 30 rounds of model validations.
-We can see that replacing GPT-4 with Llama2 results in a slight degradation of the performance of recommended models in both settings.
-This degradation is more consistent after the refinement of the model design.
-This suggests that GPT-4 is superior at proposing a reasonable next step from the trajectory and extra knowledge (optimization tasks), consistent with the existing study [53].
+4.5. Case Studies
+Different LLMs.
+To study the effect of different LLMs as meta-controllers, we use an open-source Llama2-13b to replace the GPT-4 meta-controller.
+Table 5 shows the comparative results on the initial model proposal and refinement.
+We observed that replacing GPT-4 with Llama2 results in a slight degradation of the performance of recommended models in both settings, and is more consistent in model refinement.
+This suggests that GPT-4 is superior at proposing a reasonable next step from the trajectory and extra knowledge, consistent with the existing study (Yang et al., 2023).
 Notably, because the initial model proposal phrase relies more on the quality of our elicited knowledge, Llama2 can also achieve close performance to GPT-4.
-D.4 Case Studies: Prompt Design Impact
-We studied whether fine-tuning in prompt engineering has a positive or negative effect on our work.
-Since we did not adopt any prompt fine-tuning approach, we added an alternative experiment to compare whether there is any consistent difference in the performance of LLM-based baseline GPT4GNAS [43] between using our prompt design and its original design.
-This includes adding topological graph descriptions and replacing model space descriptions and general instructions.
-No knowledge-related design from our method is being transferred because it is our key novelty.
-After prompt replacement, GHGNAS [12] became the same as GPT4GNAS [43], so we only reported one of them.
-Table 19 below shows the comparative results (averaged over five runs) after 30 rounds of refinements.
-We can see that after replacing the prompt design of the baseline with our design, the impact on the quality of the recommended models is inconsistent.
-This study justifies that our method does not improve performance by fine-tuning the prompt; instead, the main source of our performance improvement arises from our tailored model design knowledge.
-Figure 8: Short-run performance trajectories of DesiGNN compared to all the automated baselines after validating 1-30 proposals.
-Figure 9: Short-run performance trajectories of DesiGNN compared to all the automated baselines after validating 1-30 proposals.
-Table 20: Example LLM Response on Recommending Model for PubMed dataset
-Section Response (w/ Descriptional Only) Response (w/ Properties Only) Response (w/ Knowledge)
-Decision For the unseen dataset: (Architecture: [0, 0, 1, 3], Operations: [’gcn’, ’gat’, ’sage’, ’gcn’]) For the unseen dataset: (Architecture: [0, 0, 1, 3], Operations: [’gcn’, ’gat’, ’sage’, ’gin’]) For the unseen dataset: (Architecture: [0, 0, 0, 1], Operations: [’gcn’, ’gat’, ’arma’, ’graph’])
-Reason 1.
-Architecture Choice ([0, 0, 1, 3]): …… 2.
-Operations Choice ([’gcn’, ’gat’, ’sage’, ’gcn’]): - First Operation - GCN (’gcn’): The Graph Convolutional Network (GCN) is effective for aggregating neighborhood information, which is essential in a citation graph where the connectivity pattern (citations) plays a crucial role in defining the node (paper) features.
-- Second Operation - GAT (’gat’): The Graph Attention Network (GAT) introduces an attention mechanism, allowing the model to focus on the most relevant parts of the input graph structure.
-This is particularly useful in citation graphs for emphasizing influential papers.
-- Third Operation - GraphSAGE (’sage’): GraphSAGE is adept at sampling and aggregating features from a node’s neighborhood, which can help in efficiently handling the local neighborhood structures and diversity in node degrees typical in citation graphs.
-- Fourth Operation - GCN (’gcn’): Another layer of GCN helps to further refine the feature representation by aggregating higher-order neighborhood information, which can be crucial for capturing the broader context in the graph.
-This combination and sequence of operations are designed to leverage both the local and global structural information in the graph, which is critical for a citation graph where both direct and indirect citations (connections) can influence the classification of a paper into topics.
-The use of both GCN and GAT helps in effectively capturing different levels of granularity in neighbor interactions, while GraphSAGE allows for efficient computation by focusing on a sampled subset of neighbors.
-The architecture’s mix of parallel and sequential processing ensures a comprehensive feature transformation pathway, enhancing the model’s ability to learn complex patterns in the data.
-1.
-Architecture Choice ([0, 0, 1, 3]): …… 2.
-Operations Choice ([’gcn’, ’gat’, ’sage’, ’gin’]): - ’gcn’ (Graph Convolutional Network layer): Effective for capturing neighborhood information through a simple yet powerful convolutional approach.
-Given the average degree and edge count, ’gcn’ can efficiently aggregate local neighborhood features without being too computationally expensive.
-- ’gat’ (Graph Attention Network layer): Utilizes attention mechanisms to weigh the importance of nodes’ neighbors.
-This is particularly useful in graphs where nodes vary significantly in their roles and influence, as indicated by the centrality measures.
-- ’sage’ (GraphSAGE): Samples and aggregates features from a node’s neighborhood, which can handle the graph’s density and scale effectively.
-This method is adaptive and can learn from different-sized neighborhoods, which is beneficial given the variability in node connectivity.
-- ’gin’ (Graph Isomorphism Network layer): Designed to capture more nuanced graph structural information, which can be crucial for understanding deeper patterns in the graph, especially given the non-trivial graph diameter and clustering coefficient.
-This combination and sequence of operations are designed to first broadly and then selectively refine the feature representation of the nodes, considering both local and global structural properties of the graph.
-This should provide a robust model capable of handling the complexities and variabilities in the graph metrics provided.
-1.
-Architecture Choice: …… 2.
-Operation Selection: …… - ’arma’ (Auto-Regressive Moving Average): ARMA layers have been chosen in various top models, suggesting their utility in capturing more complex dependencies in graph data than standard convolutional layers.
-- ’graph’ (k-GNN): The inclusion of a k-GNN operation in the final step can extend the model’s capability to capture higher-order interactions, which might be crucial for the specific characteristics of the unseen dataset.
-3.
-Combination of Operations: The combination of GCN, GAT, and ARMA before a final GRAPH operation provides a robust sequence of transformations.
-Starting with GCN allows for a broad capture of neighborhood information, followed by the selective attention mechanism of GAT and the complex filtering of ARMA, culminating in a k-GNN layer that can capture higher-order dependencies.
-This sequence is designed to progressively refine and enhance the feature representations.
-4.
-General Observations from Top Models: The selected operations and architecture are among the most recurrent in the top-performing models from similar datasets.
-This suggests that they are likely to be effective in handling the types of graph-structured data encountered in these benchmarks, making them a prudent choice for the unseen dataset.
-By synthesizing these observations, the recommended model design aims to balance complexity and depth of feature extraction and transformation, which is crucial for effectively processing graph-structured data.
-This design is expected to perform well on the unseen dataset by leveraging proven elements from similar successful models.
-Perf. 75.40 71.56 76.60
-Table 21: Example LLM Response on Dataset Comparison for PubMed dataset
-Benchmark w/ Descriptional Inputs w/ Confident Graph Properties
-Cora Similarity Score: 0.8 Reason: Both the unseen dataset and Cora are citation graphs with papers as vertices and citations as links.
-Both use bag-of-words features and have ground-truth topics as labels.
-The main differences are in the number of features and classes, with Cora having more features and classes.
-Similarity Score: 0.1 Reason: The unseen dataset has a significantly higher graph diameter, average closeness centrality, and average betweenness centrality compared to Cora.
-Additionally, the general graph metrics such as edge count, average degree, and density are quite different, indicating a low similarity.
-Citeseer Similarity Score: 0.75 Reason: Similar to Cora, Citeseer is also a citation graph with similar characteristics to the unseen dataset.
-The differences are more pronounced in the number of features and classes, with Citeseer having significantly more features.
-Similarity Score: 0.6 Reason: Citeseer has a closer match in terms of graph diameter and average closeness centrality with the unseen dataset.
-The average betweenness centrality and general graph metrics like edge count and average degree also show some similarities, suggesting a moderate level of task similarity.
-CS Similarity Score: 0.2 Reason: CS is a co-authorship graph, differing significantly in the type of vertices (authors vs.
-papers) and the nature of features (keywords vs.
-bag-of-words).
-The task and evaluation metric are similar, but the structural differences lower the similarity score.
-Similarity Score: 0.3 Reason: CS shares a similar graph diameter with the unseen dataset but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a moderate to low similarity.
-Physics Similarity Score: 0.2 Reason: Similar to CS, Physics is a co-authorship graph with authors as vertices and co-author relationships as links, which is structurally different from the unseen dataset’s citation graph nature.
-Similarity Score: 0.2 Reason: Physics has a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and other general graph metrics like edge count and average degree, suggesting a low similarity.
-Photo Similarity Score: 0.3 Reason: Photo is a co-purchase graph, which is different in nature from a citation graph.
-However, both use bag-of-words features and have a similar number of features, slightly increasing the similarity score.
-Similarity Score: 0.1 Reason: Photo, while having a similar graph diameter, differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a low similarity.
-Computers Similarity Score: 0.3 Reason: Similar to Photo, Computers is also a co-purchase graph.
-The similarity is slightly higher due to a closer number of features and classes compared to the unseen dataset.
-Similarity Score: 0.1 Reason: Computers has a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics like edge count and average degree, indicating a low similarity.
-ogbn-arxiv Similarity Score: 0.7 Reason: ogbn-arxiv is a citation graph similar to the unseen dataset, but it uses word embeddings for features instead of bag-of-words, and has a significantly different number of classes.
-Similarity Score: 0.2 Reason: ogbn-arxiv shares a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, suggesting a low similarity.
-ogbn-proteins Similarity Score: -0.5 Reason: ogbn-proteins is structurally and fundamentally different from the unseen dataset, being a protein association graph with a multi-label binary classification task and ROC-AUC as the evaluation metric.
-Similarity Score: 0.1 Reason: ogbn-proteins has a different graph diameter and significantly different average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a very low similarity.
+Lack of Prior Knowledge in LLMs.
+The inherent knowledge gap illustrated in Figure 1(a) refers to the lack of prior knowledge in LLMs about the top-performing models of benchmark datasets within the NAS-Bench-Graph (Qin et al., 2022) model space.
+We analyze LLM responses in which the descriptive inputs or graph properties were directly sent to the LLMs for model suggestions.
+The results demonstrate that LLMs are only able to suggest commonplace layer connections and operations based on user-provided descriptive inputs or key graph properties.
+We have tested all benchmark datasets in NAS-Bench-Graph and found that only two macro lists and three operation lists would be recommended by LLMs regardless of the datasets: (1) Architecture: [0, 1, 2, 3] and [0, 0, 1, 3]; (2) Operations: [’gcn’, ’gat’, ’sage’, ’skip’], [’gcn’, ’gat’, ’sage’, ’gcn’], and [’gcn’, ’gat’, ’gin’, ’sage’].
+After providing tailored knowledge derived from our framework, the LLM-recommended model becomes specialized and shows significant performance improvements.
+Artificial Hallucination in LLMs.
+In Figure 1(b), we examine the phenomenon of “artificial hallucination,” or the external noise issue, in LLMs when comparing the similarities between unseen and benchmark datasets (Section 3.2).
+Figure 6 (leftmost) illustrates that the empirically most similar benchmark datasets to PubMed are CS, Physics, and Citeseer, in descending order of similarity.
+When employing our Graph Understanding method in Section 3.1, which leverages confident graph properties, the top three most similar datasets are Citeseer, CS, and Physics (align with the empirical ground truth).
+However, the current practice that relies solely on descriptive inputs identifies Cora, Citeseer, and ogbn-arxiv as the top three, which does not align with the empirical ground truth.
+This discrepancy arises because LLMs overly rely on the shared characteristic of being citation graphs, assuming that citation datasets like PubMed, Cora, Citeseer, and ogbn-arxiv should have similar model design preferences.
+This reveals that relying solely on external descriptive inputs is insufficient to capture the similarities between datasets, which can overwhelm other pertinent information, leading to inaccurate task association and knowledge retrieval.
+5. Conclusion
+We present DesiGNN, a knowledge-centered framework that demonstrates deep expertise in proficiently designing data-aware GNNs and eliminates external noise by systematically converting accumulated model design experience into fine-grained meta-knowledge well suited to meta-learning with LLMs.
+Leveraging a performance-oriented task similarity grounded in empirically validated graph properties and adaptive knowledge elicitation via LLMs, DesiGNN bridges the gap between literature insights and practical, data-specific model deployment.
+DesiGNN integrates graph understanding, effective knowledge retrieval, and strategic model suggestion and refinement under a Bayesian perspective, enabling tailored GNN designs for diverse graphs in a single shot or with a few attempts.
+Extensive experiments demonstrate that DesiGNN expedites the design process by delivering top-tier GNN designs within seconds and achieving superior performance at minimal cost.
+Acknowledgements.
+Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No.
+2023YFF0725100, National Science Foundation of China (NSFC) under Grant No.
+U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No.
+2024A0505040012, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Key Areas Special Project of Guangdong Provincial Universities 2024ZDZX1006, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant, HKUST-Webank joint research lab, 2023 HKUST Shenzhen-Hong Kong Collaborative Innovation Institute Green Sustainability Special Fund, from Shui On Xintiandi and the InnoSpace GBA, and HKUST(GZ) - CMCC(Guangzhou Branch) Metaverse Joint Innovation Lab under Grant No.
+P00659.
+Jiachuan Wang’s work is supported in part by JST CREST (JPMJCR22M2).
+Shimin Di’s work is supported by National Science Foundation of China (NSFC) under Grant No.
+62506075.